### v1

# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

Abstract
Steering vision foundation models at inference time without retraining or access to large labeled datasets is a desirable yet challenging objective, particularly in dynamic or resource-constrained settings. In this paper, we introduce Visual Sparse Steering (VS2), a lightweight, test-time method that guides vision models using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 surpasses zero-shot CLIP by 4.12% on CIFAR-100, 1.08% on CUB-200, and 1.84% on Tiny-ImageNet. We further propose VS2++, a retrieval-augmented variant that selectively amplifies relevant sparse features using pseudo-labeled neighbors at inference time. With oracle positive/negative sets, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to better align the sparse features learned through the SAE reconstruction task with those relevant for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS). By incorporating a prototype-alignment loss during SAE training, using labels only during training while remaining fully test-time unsupervised, PASS consistently, though modestly, outperforms VS2, achieving a 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
1 Introduction
Foundation Models (FMs) have exhibited strong generalization capabilities across a wide range of domains, from classification tasks to open-ended generation  clip ; sam ; llama2 ; llava ; gpt4 . Despite their impressive performance, these models largely function as black boxes, restricting their controllability, and reliability in practice. The standard approach to edit a model’s behavior involves first collecting a labeled dataset resembling the target objective, followed by a supervised fine-tuning (SFT) procedure upon the holistic zhang2023instruction ; dong2023abilities or partial (e.g., PEFT lester2021power ; lora ; ia3 ) parameter space. Nevertheless, this procedure is cumbersome and can become prohibitively expensive given the model scale. Consequently, Steering Vector (SV) methods turner2023activation ; zou2023representation ; liu2024incontext , as an alternative, have gained popularity in the Large Language Model (LLM) domain as they can effectively modulate the model’s generation at inference-time in a training-free manner. While SVs have shown promise in LLMs, their functionality in the vision domain remains unclear.
In this work, we investigate the functionality of steering vectors in vision foundation models (VFMs). In particular, we use zero-shot classification tasks as a testbed to explore how SVs can be leveraged to upweight relevant visual features. Conventional SVs operate by first constructing a directional vector within the latent space to indicate a desired change of the functionality. It then linearly combines the directional vector with the original latent embeddings to craft output embeddings. In LLMs, SVs rely on two symbolic anchors that are inherently interpretable to determine the functional direction of the SV. For example, a happiness SV can be extracted using the latent embedding arithmetic corresponding to Happy - Sad. The formulation of SVs in LLM poses two critical challenges in vision domain. Given an image at inference-time: (1) Unlike symbolic language tokens, there exists tremendous redundancy in visual representations; consequently, it is unclear what features should be steered; (2) Without inherently interpretable anchors (e.g., happy, sad), towards which direction to steer is not obvious.
(a) VS2: Visual Sparse Steering w/o Contrastive Data
(b) VS2++: Selective Visual Sparse Steering
Figure 1: Overview of VS2 and VS2++. At inference, VS2 finds sparse latent concepts, and steers the original embedding towards the direction of amplifying those sparse features. When additional caching data is available, VS2 selectively enhances certain features while suppressing others in a contrastive manner.
To address the aforementioned challenges, we propose Visual Sparse Steering (VS2). Drawing inspiration from the interpretability provided by sparse auto-encoders (SAEs) hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , we hypothesize that the sparse features in a SAE’s latent space can capture key aspects of a visual embedding. VS2 upweights these features equally to construct a steering vector for classification. This equal upweighting of features is imprecise in that we expect different features to be more relevant than others for classification tasks. Despite this, we show that the simple equal upweighting of features still improves zero-shot classification.
To construct a more precise steering vector, we propose VS2++, which leverages additional unlabeled data to construct steering vectors for classification. For a given target image, VS2++ retrieves the Top-N most similar visual embeddings, and constructs positive and negative groups according to generated pseudo labels [^1] In practice, pseudo labels can be gathered using either unsupervised clustering methods or zero-shot classification results. The VS2++ SV is then constructed via subtracting the average SAE features of the negative group from the positive group. Intuitively, VS2++ upweights the SAE features most relevant for distinguishing the positive and negative groups. In summary, VS2 and VS2++ aim to enhance downstream classification performance while retaining the interpretability of SAEs. To further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS).
VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP  radford2021learning by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively, without any external data or supervision. When using external unlabeled image data, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Finally, by adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Our main contributions are:
  • We investigate the functionality of steering vectors in vision domain, and propose VS2 that leverages sparse feature learning to construct steering vectors for zero-shot classification.
  • We further propose VS2++ for scenarios where additional unlabeled data is available to selectively amplify important concepts. In this case, we use VS2 on pseudo-positive and -negative data to upweight relevant features for classification.
  • We propose Prototype-Aligned Sparse Steering (PASS), a lightweight method that introduces a prototype-alignment loss during SAE training using class labels only during training while remaining fully test-time unsupervised.
  • We empirically demonstrate that VS2, VS2++, and PASS can leverage SAEs to substantially improve zero-shot classification performance, outperforming baseline methods by a large margin under their corresponding regimes, while also retaining interpretability.
2 Related Work
Mechanistic Interpretability and Sparse Autoencoders.
Several traditional approaches exist for interpretability in vision models, including feature visualizationsimonyan2014visualising ; zeiler2014visualizing ; olah2017feature and network dissectionbau2017network ; oikarinen2022clip . Mechanistic interpretability seeks to systematically analyze and understand neural networks elhage2021mathematical ; olah2020zoom , but it faces challenges due to polysemantic neurons i.e. units that activate in response to multiple, seemingly unrelated inputs elhage2022toy . This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations elhage2022toy . Sparse Autoencoders (SAEs) have been shown to mitigate superposition by applying sparse dictionary learning to model internals sharkey2022taking ; bricken2023monosemanticity . Recent efforts have leveraged SAEs to uncover interpretable features within LLMs, revealing latent units tied to grammar rules, style patterns, and factual knowledge templeton2024scaling ; cunningham2023sparse ; gao2024scaling .  joshi2025identifiable refines this by training autoencoders on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision. While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
While some studies investigate SAEs for exploratory feature analysis or generative modeling in the visual domain hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , they primarily address concept discovery rather than performance enhancement. Other attempts in the vision domain try to create new SAE architectures for concept extraction  fel2025archetypal ; thasarathan2025universal . In contrast, our method leverages SAEs not only to interpret but also to actively steer vision models, yielding classification improvements alongside interpretable latent representations.
Similar to our line of work, joseph2025steering validate CLIP’s steerability and its efficacy in suppressing typographic attacks. Patch-SAElim2024sparse improves zero-shot classification by applying latent masking based on class-level activations. Specifically, it aggregates SAE activations from the training split to identify the top-$k$ most frequently activated latents per class, which are then used to mask active latents at test time. Our approach differs in a key aspect: we do not rely on activation aggregations from the same class in training but our method is instead fully test-time unsupervised. This design avoids reliance on class-conditioned statistics, making our approach more flexible and applicable to a wide range of tasks without requiring dataset-specific activation distributions. Specifically, our method (i) operates entirely in the vision latent space, (ii) requires no class labels, external vocabularies, or gradient updates, and (iii) improves zero-shot classification by amplifying sparse SAE features rather than replacing or suppressing activations.
Steering Vectors.
Steering vector (SV) methods turner2023activation ; park2023the ; hernandez2024linearity ; NIPS2013_9aa42b31 , also known as representation engineering zou2023representation , construct a directional vector (or task vector) and apply it in the latent space to change the target model’s behavior at inference-time. SVs have gained popularity in LLM/MLLM domain as it can improve generation security liu2024incontext , promoting truthfulness NEURIPS2023_81b83900 , mitigating hallucination vista , as well as improving inference efficiency liu2024incontext ; i2cl . Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts gavrikov2025can . In contrast, we differentiate from this line of work by focusing on Vision Foundation Models directly steering representations in the latent space without any external language signal.
Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions bayat2025steering ; makelov2024sparse . Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled concept pairs, making it more adaptable to general visual representations. Furthermore, its applicability to vision foundation models (VFM) remains unclear. The proposed visual sparse steering (VS2) method demonstrates that SV methods, combined with sparse autoencoders, can be utilized to enhance image features, improving zero-shot classification.
3 Method
3.1 Preliminaries
Sparse Autoencoder (SAE). A SAE consists of an encoder-decoder pair that maps an input vector ${\bm{x}}$ into a latent representation ${\bm{z}}$, then reconstructs ${\bm{x}}$ from ${\bm{z}}$. Formally, we define the encoder as:
${\bm{z}}=\mathrm{ReLU}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}\Bigr{)},$ (1)
where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input. The decoder then reconstructs ${\bm{x}}$ as:
$\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}},$ (2)
where ${\bm{W}}_{\mathrm{dec}}\in{\mathbb{R}}^{d\times n}$ is the decoder weight matrix. To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective is:
$\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ (3)
In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  gao2024scaling explicitly enforces the sparsity by using a top-$k$ selection mechanism:
${\bm{z}}=\mathrm{TopK}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})\Bigr{)},$ (4)
where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest. This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
3.2 Visual Sparse Steering
In this section, we introduce Visual Sparse Steering (VS2), a novel framework that leverages sparse features to steer neural representations toward meaningful semantic directions. Our approach targets two interconnected objectives, i.e., (1) explicitly identifying semantically meaningful concepts embedded within latent representations and (2) utilizing these interpretable concepts to define steering directions that enhance downstream task performance by emphasizing relevant semantic information and reducing spurious correlations. Unlike standard steering methods that rely on explicitly curated positive and negative examples, our approach utilizes Sparse Autoencoders (SAEs) to automatically discover and isolate sparse concepts that assist in defining a steering direction. By selectively amplifying these concepts within the latent space, VS2 generates conceptually grounded steering vectors that improve model interpretability, enhance classification accuracy, and robustly mitigate reliance on superficial or irrelevant cues.
3.2.1 VS2: Visual Sparse Steering
Learning latent concept space. Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding ${\bm{x}}$ into a sparse concept activation vector:
${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|}).$ (5)
By imposing sparsity constraints described in Sec. 3.1 on the concept activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful concepts. To map the concepts into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these concepts:
$\tilde{{\bm{x}}}=D_{c}({\bm{c}}).$ (6)
Since explicit supervision for concept identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful concepts.
Constructing steering vector. At inference stage, emerged sparse concepts are deemed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$. In implementation, we simply scale the original concept activations ${\bm{c}}$ by a factor of $\gamma$ to amplify their effects:
${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}.$ (7)
Reconstructing this modified concept vector ${\bm{c}}^{\prime}$ provides a conceptually steered embedding:
$\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime}).$ (8)
Since reconstruction steps inherently introduce approximation errors, we define the steering vector ${\bm{v}}$ as:
${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}},$ (9)
which explicitly represents the semantic shift induced by upweighting and mitigates the effects of the reconstruction errors.
Steer the target emebedding. Obtaining the steering vector ${\bm{v}}$ that aims to amplify the latent concepts embedded in ${\bm{x}}$, we shift the original embedding by adding the steering vector back as:
$\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}.$ (10)
where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale:
$\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ (11)
Intuitively, this procedure precisely moves the embedding ${\bm{x}}$ along directions defined by semantically relevant, interpretable concepts, thus improving interpretability and enhancing robustness against spurious correlations. We refer Fig. 1 (upper) for a schematic overview of VS2.
3.2.2 VS2++: Visual Sparse Steering with Cached Data
In VS2, we leverage the learned SAEs to implicitly decide the critical concepts for a given latent embedding, and treat all identified concepts equally important. One natural speculation is that there are certain concepts that are unique and more important to the target underlying distribution, while others are more general even spuriously correlated. To this end, we propose VS2++ that allows the access of cached training embedding at test-time to enable a more nuanced control over the concept steering.
Constructing Contrastive Groups. Given a query embedding ${\bm{x}}_{q}$ and a cached dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
$\mathcal{N}_{N}({\bm{x}}_{q})\;=\;\underset{\begin{subarray}{c}\{{\bm{x}}_{i}\}\subseteq\{{\bm{x}}_{i}\}_{i=1}^{M}\\ |\mathcal{N}_{N}({\bm{x}})|=N\end{subarray}}{\mathrm{arg\,max}}\;S({\bm{x}}_{q},{\bm{x}}_{i}),$ (12)
where $S(,)$ denotes the measurement of cosine similarity.
After retrieving $\mathcal{N}_{N}({\bm{x}}_{q})$, we obtain the pseudo-label for the query embedding $\hat{y}_{q}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{q})$, where $P(y\,\mid\,{\bm{x}})$ is the classifier’s predicted probability for label $y$. Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$.
We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding:
${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ (13)
Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors:
${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$ (14)
Selective Steering. For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first applies VS2 (Eq. 7,8,9) to acquire a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying concepts within the given embedding. We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$. We then average these per-embedding steering vectors within each group, yielding the positive anchor
$\bar{{\bm{v}}}^{p}\;=\;\frac{1}{\bigl{|}{\mathbb{S}}^{+}({\bm{x}}_{q})\bigr{|}}\,\sum_{{\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})}{\bm{v}}_{i}^{p},$ (15)
and the negative anchor
$\bar{{\bm{v}}}^{n}\;=\;\frac{1}{\bigl{|}{\mathbb{S}}^{-}({\bm{x}}_{q})\bigr{|}}\,\sum_{{\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})}{\bm{v}}_{j}^{n}.$ (16)
Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor:
${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ (17)
and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
$\hat{{\bm{x}}}_{q}\;=\;{\bm{x}}_{q}\;+\;\lambda\,{\bm{v}},\quad\hat{{\bm{x}}}_{q}\;=\;\frac{\hat{{\bm{x}}}_{q}\,\cdot\,\|{\bm{x}}_{q}\|_{2}}{\|\hat{{\bm{x}}}_{q}\|_{2}}.$ (18)
This selective scheme thus anchors the steering directions in concepts consistently activated within the positive group, while suppressing undesired features highlighted by the negative group. Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases.
4 Experiments
Table 1: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and Without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Zero-shot (no retrieval) $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 ) RAG-enhanced (oracle cache) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle cache) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
Experimental Setup.
We test VS2 and VS2++ on CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny , and CUB-200 wah2011caltech three datasets, covering standard, complex, and fine-grained classification tasks. These datasets vary in scale, diversity, and class distribution, allowing us to assess VS2 and VS2++ in both general and imbalanced classification scenarios. For vision foundation model, we use CLIP radford2021learning with ViT-B/32, and ViT-B/16 backbones. We train top-$k$ Sparse Autoencoders (SAEs) on the CLS token embeddings extracted from every layer of CLIP’s vision transformer, using the training data of each dataset to learn sparse latent representations. More details regarding the training of the top-$k$ Sparse Autoencoders can be found in the Appendix C.
4.1 Visual Sparse Steering methods
4.1.1 VS2: Visual Sparse Steering without Contrastive Data
Baselines.
We evaluate the effectiveness of VS2 in the absence of external contrastive data. As baselines, we report: (1) the zero-shot performance of the original CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{CLIP}_{\text{REC}}^{\text{F}}$, which uses the SAE-reconstructed CLS token from the final layer without any steering; and (3) $\text{CLIP}_{\text{REC}}^{\text{A}}$, which uses the reconstructed CLS tokens from all layers. These baselines isolate the effect of reconstruction alone, without applying any modification or steering vector. All top-$k$ SAEs are trained on CLS token representations across all transformer layers. In addition to these baselines, we report $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ that applies steering by upweighting the top-$k$ sparse features of the reconstructed CLS token of the final layer with a scaling factor of $\gamma=1.5$. This variant isolates the impact of steering the SAE’s latent activations by amplifying their top-$k$ before the CLS embedding is reconstructed, letting us quantify the benefit attributable to this targeted up-weighting alone. In Appendix G, we provide the pseudocode for main differences between the proposed baselines and VS2. In Appendix H, we also provide comparative results of our method compared to other baselines such as Splice (bhalla2024interpreting, ). Note that Splice uses an external vocabulary to construct latent features; in contrast, we use unsupervised SAEs to learn latent features for VS2 and VS2++.
Results.
From Table 1, we demonstrate that using only SAE reconstructions generally reduces zero-shot performance in all datasets; we hypothesize this is due to reconstruction errors in SAEs (engels2025decomposing, ). However, when SAE features are amplified in the reconstruction, performance can improve, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ for CIFAR-100. Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% using the ViT-B/32 and ViT-B/16 backbones respectively. However, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ is worse than the CLIP baseline on CUB-200 and Tiny-ImageNet. In contrast, our method VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. In summary, we demonstrate that our proposed steering vectors using SAEs can improve downstream classification performance on vision tasks without relying on any external data or supervision. In Appendix A, we also provide evidence that the sparse features are interpretable; they can capture subtle class-specific attributes and concepts such as birds with gray upperpart, white underpart, white-colored throat and black eyes. Finally, in Appendix B, we provide a sensitivity analysis of VS2 hyperparameters $\gamma$ and $\lambda$ showing the robustness of our proposed methods.
4.1.2 VS2++: Visual Sparse Steering with Selectively Amplifying Features
In VS2, we leverage learned SAEs to identify critical concepts from a latent embedding, treating all activated concepts as equally important. However, some concepts may be more discriminative or unique to the target class, while others may be generic or even spuriously correlated with a class label. To enable finer control over concept importance, we introduce VS2++, which accesses cached training data at test time to guide steering more selectively. As a proxy for identifying salient features, we adopt a Retrieval-Augmented Generation (RAG) approach: we retrieve the top-$50$ most similar training images to the test input using DINOv2 oquab2023dinov2 and average their sparse activations. To construct the positive and negative sets as shown in equations 13, 14, we assign pseudolabels using a CLIP classifier; images with the most frequent pseudolabel form the positive set, while the remaining images serve as negatives. This setup corresponds to the Non-oracle cache in Table 1. For comparison, we also report an unrealistic scenario that serves as an upper-bound i.e. Oracle cache setting that uses ground-truth labels to define positives and negatives.
Baselines.
As a baseline in the Retrieval-Augmented Generation (RAG) pipeline lewis2020retrieval , we combine the test query embedding with a weighted aggregation of retrieved document embeddings. Additional details on the weighting strategy are provided in Appendix D. We also include a non-SAE variant that constructs a steering vector by averaging the CLIP embeddings of the positive and negative sets and taking their difference, yielding a purely contrastive direction without relying on Sparse Autoencoders.
Results.
In Table 1, we present the results for VS2 ++ when external data is available under both RAG-enhanced settings where we have oracle and non-oracle cache data. When we have oracle positive and negative sets, we observe that VS2++ surpasses all other methods under almost every setting. Weighted RAG performs worse than the steering vector approaches (both CLIP and VS2++), suggesting that directionality is more informative than proximity. In the more realistic scenario where we do not have oracle positive and negative sets, we observe drops in accuracies. We hypothesize that this degradation is due to inaccuracies in pseudolabeling, resulting in the amplification of spurious directions in the SAE latent space. However, across the six configurations, VS2++ outperforms the CLIP Steering Vector in four cases and falls behind only in the ViT-B/32 runs for CIFAR-100 and Tiny-ImageNet (with CIFAR-100 essentially on par). These results suggest that there remains significant room for improvement in actively selecting which sparse features to steer, especially under noisy or weakly supervised retrieval scenarios. In Appendix I, we conduct ablation studies on the top-$N$ retrieved images that indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
4.1.3 Fine-Grained Per-Class Accuracy Analysis
To understand where Visual Sparse Steering helps most, we compute per-class top-1 accuracy on CIFAR-100 with a ViT-B/32 backbone and report the ten largest improvements in Table 2. The results reveal a long-tailed pattern: VS2 lifts accuracy by up to 25 % in individual classes (e.g. lawn-mower $\rightarrow$ tractor, pine-tree $\rightarrow$ forest), while VS2++ reaches gains of 38 % when high-quality neighbours are available (e.g. bee $\rightarrow$ spider, lion $\rightarrow$ tiger, whale $\rightarrow$ flatfish). In all cases, the corrected predictions involve visually or taxonomically proximate categories, indicating that sparse steering primarily disambiguates fine-grained confusions rather than providing a uniform boost. This behaviour aligns with our qualitative findings in the Appendix A.2, where the top-$k$ SAE features capture subtle class-specific attributes (birds with gray upperpart, white underpart, white-colored throat and black eyes etc.) that are otherwise entangled in the original CLIP embedding. A more detailed list of per-class performance accuracies of VS2 and VS2++ is provided in Table 9 in the Appendix.
Table 2: Top-10 class gains on CIFAR-100. Numbers are top-1 accuracies; green = absolute gain over zero-shot (ZS).
Class ZS VS2 (+$\Delta$) Mislabel Why visually confusing tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower Both are wheeled agricultural machines; similar viewed side-on. forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree Dense tree canopies often dominated by tall conifers. man $53.0$ 74.0 $\uparrow$(+21.0) boy Male human silhouette; age difference subtle at CIFAR resolution. bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck Rectangular vehicle profile with windows and wheels. snake $61.0$ 76.0 $\uparrow$(+15.0) worm Long, legless bodies with smooth textures. woman $57.0$ 71.0 $\uparrow$(+14.0) girl Female human figures; height/size cues are small. trout $18.0$ 32.0 $\uparrow$(+14.0) aquarium_fish Similar fish shape and reflective scales in water scenes. cockroach $49.0$ 62.0 $\uparrow$(+13.0) beetle Small, dark exoskeleton insects with segmented bodies. bridge $72.0$ 83.0 $\uparrow$(+11.0) sea Large spans photographed above wide blue water areas. rose $58.0$ 68.0 $\uparrow$(+10.0) poppy Red, multi-petal flowers with central dark core.
(a) VS2
Class ZS VS2++ (+$\Delta$) Mislabel Why visually confusing spider $53.0$ 91.0 $\uparrow$(+38.0) bee Small multi-legged insects; dark radial silhouettes. tiger $48.0$ 84.0 $\uparrow$(+36.0) lion Large orange/brown big cats; stripes vs. mane often unclear at low resolution. flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale Flattened marine shapes against blue water backgrounds. possum $30.0$ 65.0 $\uparrow$(+35.0) hamster Small brown furry mammals with rounded bodies. lizard $39.0$ 74.0 $\uparrow$(+35.0) snake Slender reptiles; legs on lizard hard to spot in $32\times 32$ images. sweet_pepper $58.0$ 92.0 $\uparrow$(+34.0) orange Round orange-coloured produce with glossy skins. raccoon $49.0$ 83.0 $\uparrow$(+34.0) skunk Mid-sized nocturnal mammals with dark fur and contrasting markings. shrew $46.0$ 79.0 $\uparrow$(+33.0) skunk Small ground mammals; dark elongated silhouettes blur distinguishing cues. snail $53.0$ 86.0 $\uparrow$(+33.0) mushroom Brown dome (shell vs. cap) resting on the forest floor. beetle $49.0$ 82.0 $\uparrow$(+33.0) cockroach Dark, shiny exoskeletal insects with similar body shapes.
(b) VS2++
4.2 Prototype-Aware Sparse Steering Vectors
Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE). These directions correspond to latent concepts that, ideally, align with class-discriminative features. Steering in these directions upweights what the model has learned to be important during reconstruction. This raises a central hypothesis: the reconstruction task itself is to some extent sufficient to uncover features that are also relevant for classification. In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task. In this section, drawing inspiration from prototype theory rosch1973natural , we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
The limited improvements observed on fine-grained datasets like CUB-200 and Tiny-ImageNet suggest that the challenge lies not just in identifying sparse features, but in uncovering the correct ones. This shifts the central question from “What are the most important sparse features to select?” to a deeper inquiry: “Can the reconstruction objective alone reliably capture features that are most useful for classification and if not, how can task-relevant information be effectively incorporated?”
Oracle steering with known prototypes.
Following prototype theory, we collect for every class the ten images that ViT-B/32 CLIP classifies with the highest confidence; these serve as oracle prototypes. Given the true class label $y$, we build a prototype steering vector by averaging their latent sparse features. VS2 using oracle prototypes can lift CLIP to 97.5% on CIFAR-100, 91.04% on CUB-200, and 90.1% on Tiny-ImageNet. This confirms the existence of discriminative sparse directions. In Appendix E we further examine the discriminative ability (measured by orthogonality) of these steering vectors. Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
Prototype-aligned SAE (PASS).
Above, we constructed steering vectors using the pretrained SAE features of oracle prototypes. Now, we consider whether prototypes can be used to learn new, more informative SAE features. We assume during SAE training that we have access to class labels. Then, to the SAE loss we add a regularization term which encourages SAE features to be close to their class mean. That is, for a training sample $i$ with latent sparse feature  $\mathbf{z}_{i}$ and class mean $\bar{\mathbf{z}}_{\text{class}(i)}$ we minimize
$\mathcal{L}=\mathcal{L}_{\text{recon}}+w_{\text{aux}}\;\bigl{\lVert}\mathbf{z}_{i}-\bar{\mathbf{z}}_{\text{class}(i)}\bigr{\rVert}_{2}^{2},$ (19)
where $w_{\text{aux}}$ controls the strength of the prototype-alignment term relative to the reconstruction loss and is set to 0.8. We refer to the resulting steering method as PASS (Prototype-Aligned Sparse Steering). Although PASS uses class labels during SAE training, it remains fully test-time unsupervised. Empirically, in Table 1, we observe that PASS outperforms VS2 across all datasets, with particularly substantial gains on CIFAR-100. However, this improvement comes at the cost of requiring labels for each training sample during SAE training. Gains are modest on CUB-200 and Tiny-ImageNet, and we thus hypothesize that classes which share many features require richer or multi-prototype guidance which is an intriguing avenue for future work.
5 Conclusion
We introduce Visual Sparse Steering VS2 and its retrieval-augmented variant VS2++, two test-time methods that steer vision foundation models using sparse features obtained from top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. To control concept importance more precisely, we propose VS2++. With oracle positive/negative sets, VS2++ surpasses every baseline in nearly every configuration, achieving absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. When those sets are noisy, its accuracy drops but still surpasses the standard Steering Vector in most cases, highlighting the need for better feature selection under weakly supervised retrieval. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we proposed Prototype-Aligned Sparse Steering (PASS). By adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Selecting appropriate prototypes or exploring multi-prototype guidance is an intriguing avenue for future work.
Appendix
Appendix A Decoding the Sparse Latent Space: Insights from SAEs
Our proposed methods, VS2 and VS2++, achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs). We hypothesize that SAEs identify meaningful sparse concepts, which in turn guide the steering mechanisms. To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned concepts.
A.1 Quantitative Evaluation of Concept Significance.
To further evaluate the role of interpretable sparse features, we manipulate the top-$k$ most active latent dimensions extracted via Sparse Autoencoders. This experiment examines whether these features are critical for classification and how model predictions change under different modifications. We explore the following manipulation settings:
  1. 1.
    Zeroing-Out ($\gamma=0$): We set the top-$k$ most active features to zero before applying the steering vector. This removes their influence while preserving the remaining latent structure.
  2. 2.
    Negation ($\gamma=-1$): We invert the sign of the top-$k$ features before applying the steering vector, effectively pushing the representation in the opposite direction. This tests whether these dimensions encode class-discriminative information.
Table 3 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention. We observe that negating or zeroing the top-$k$ most important concepts consistently degrades performance across all datasets. This result confirms that the concepts learned by the Sparse Autoencoders are essential for classification.
Table 3: Effect of manipulating top-$k$ sparse codes. Zero-shot accuracy (%) drops sharply when dominant sparse features are zeroed ($\gamma=0$) or negated ($\gamma=-1$), confirming their importance.
Modification CIFAR-100 CUB-200 Tiny-IN
$\text{CLIP}_{\text{ZS}}$ $61.07$ $38.68$ $56.64$
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $37.30$ $39.49$
+ Zero-out ($\gamma=0$) $1.71$ $1.14$ $16.11$
+ Negate ($\gamma=-1$) $0.06$ $0.00$ $0.82$
A.2 Qualitative Evaluation of Concept Significance.
We qualitatively investigate the concepts learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE). Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature. Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes. For the qualitative evaluation of concept significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations. Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images. Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension. The intuition is that if a specific concept frequently emerges among the top-activating images for a particular dimension, that dimension is strongly associated with that concept.
For feature 511 as shown in Figure 2 (left), the activated images exhibit consistent semantic characteristics, including a gray upper part and a white underpart. Notably, this feature predominantly activates for images from similar but different classes, specifically different types of Gulls, such as Western, California, Herring, and Slaty-Backed Gulls. We observe that the top-activating images for these classes share all concepts in common which is not the case with dimension 3067. For feature 3067, as shown in Figure 2 (right) we observe that the top-activating images share common visual attributes, such as a white-colored throat and black eyes. Through human qualitative evaluation, we find that features in the sparse latent space capture meaningful visual concepts, grouping semantically similar images together, either from the same class or across different classes, as long as they share underlying conceptual similarities.
Figure 2: Concept coverage analysis of learned sparse latent features in the Sparse Autoencoder (SAE). Each subfigure illustrates the top-activating images for two different SAE dimensions, highlighting the consistency of shared visual concepts among the highest-activated examples. The analysis demonstrates how sparse features capture meaningful semantic attributes, grouping semantically similar images either within or across classes. Left: Top-activating images for feature 511. The images predominantly belong to different classes of Gulls (e.g., Western, California, Herring, and Slaty-Backed Gulls), yet they share consistent visual characteristics such as a gray upper part and a white underpart. This suggests that feature 511 captures a semantically meaningful concept spanning multiple related categories. Right: Top-activating images for feature 3067. The images share distinct visual attributes, including a white-colored throat and black eyes. However, unlike feature 511, these images belong to more diverse categories, indicating that this latent dimension captures a broader concept that generalizes across different classes.
Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
We analyze how the two key hyperparameters in VS2 (i) the sparse-feature amplification $\gamma$, and (ii) the steering vector scale $\lambda$ affect downstream accuracy. In the absence of contrastive supervision, these parameters govern how strongly we amplify sparse activations and how far the embedding is shifted in feature space. We sweep both values across a grid on three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet using ViT-B/32 backbone.
Figure 3 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees. Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy. For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$. Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse codes and embedding distortion. The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
(a) CIFAR-100
(b) CUB-200
(c) Tiny-ImageNet
Figure 3: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$. All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$. Accuracy degrades if either parameter becomes too large.
Appendix C Training top-$k$ Sparse Autoencoders
We follow CLIP [18] with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token. Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer. We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned. We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units. Training largely follows [31] and uses  [53], with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
We monitor reconstruction quality with the fraction of variance unexplained (FVU; lower is better), defined as
$\text{FVU}\;=\;\frac{\lVert\mathbf{X}-\hat{\mathbf{X}}\rVert_{F}^{2}}{\lVert\mathbf{X}-\bar{\mathbf{X}}\rVert_{F}^{2}},$ (20)
where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance). FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 00 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean. In Table 4, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
Table 4: Fraction of variance unexplained (FVU; lower is better). Each SAE has 4× expansion; sparsity $k=64$ for ViT-B/32 and $k=256$ for ViT-B/16.
Dataset ViT-B/32 (k=64) ViT-B/16 (k=256)
CIFAR-100 $0.2812$ $0.1166$
CUB-200 $0.2487$ $0.1653$
Tiny-ImageNet $0.5060$ $0.3018$
Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline. We use DINOv2 [50] to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
$\mathbf{E}=\alpha\mathbf{q}+(1-\alpha)\sum_{j=1}^{k}w_{j}\mathbf{r}_{j},$
where $\mathbf{q}$ is the query embedding, $\mathbf{r}_{j}$ the $j$-th retrieved embedding, and $w_{j}$ the normalized similarity weight:
$w_{j}=\frac{s_{j}}{\sum_{i=1}^{k}s_{i}},\quad\text{where }s_{j}=\text{sim}(\mathbf{q},\mathbf{r}_{j}).$
The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set.
We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 4 shows results for CIFAR-100, and Tiny-ImageNet.
Findings.
CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance. On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result. This suggests that the retrieval set carries strong distributional signal in cluttered datasets. Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context.
The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines. For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ work best; for noisy domains, more reliance on the retrieval pool is beneficial.
(a) CIFAR-100
(b) Tiny-ImageNet
Figure 4: RAG sensitivity to $\alpha$ and top-$k$. Accuracy varies with the weight $\alpha$ on the original query and the number $k$ of retrieved images. Larger $k$ often introduces noise; smaller $\alpha$ performs better on cluttered datasets.
Appendix E Are exemplar-derived directions really distinct?
A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B. Using the oracle prototypes as described in Section 4.2, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder. Most pairs have low similarity ($\!\text{mean}\!=\!0.23$), yet a non-negligible tail reveals strongly overlapping directions. Table 5 lists the ten highest-overlap pairs:
Table 5: Top-10 most overlapping prototype steering directions on CIFAR-100. High cosine similarity indicates that the two classes share visual attributes that the SAE encodes along nearly the same latent axis.
Rank Class 1 Class 2 Cosine $\uparrow$
1 beetle cockroach $0.91$
2 mouse shrew $0.89$
3 dolphin shark $0.84$
4 otter seal $0.84$
5 dolphin whale $0.84$
6 possum raccoon $0.84$
7 snake worm $0.83$
8 oak tree willow tree $0.83$
9 ray shark $0.81$
10 bowl cup $0.80$
These high-overlap pairs are semantically plausible confusions (e.g. beetle vs. cockroach or dolphin vs. whale), confirming that exemplar steering directions tend to align for visually or taxonomically proximate classes. In downstream applications, a simple orthogonalization step may help reduce feature overlap between sparse directions. Investigating principled ways to encourage orthogonality during SAE training is a promising direction for future work.
Appendix F Ablation Study of Expansion factor and top-$k$
In Table 6, we present the downstream task accuracy of CLIP ViT-B/32 using various values of expansion factor and $k$. Across a 4 $\times$ range in width and an 8 $\times$ range in sparsity, top-1 accuracy fluctuates by less than one percentage point evidence that VS2’s performance is largely insensitive to the precise SAE capacity–sparsity trade-off. Additionally, in Table 7, we present the average cosine similarities of different steering vectors coming from various configurations of SAEs in terms of expansion factor and top-$k$.
Table 6: VS2 accuracy as a function of SAE width (expansion factor) and sparsity (top-$k$). Numbers are top-1 / top-5 (%). The best result is boldfaced; every other configuration is within of the optimum, highlighting the method’s robustness to architectural choices.
SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
4 $\times$, $k$=128 64.61 87.95
8 $\times$, $k$=128 64.56 87.76
4 $\times$, $k$=64 64.54 87.79
16 $\times$, $k$=64 64.54 87.78
8 $\times$, $k$=64 64.52 87.96
10 $\times$, $k$=128 64.43 87.81
16 $\times$, $k$=512 64.42 87.71
8 $\times$, $k$=512 64.40 87.91
4 $\times$, $k$=256 64.34 87.62
8 $\times$, $k$=256 64.29 87.80
16 $\times$, $k$=256 64.28 87.75
4 $\times$, $k$=512 64.12 87.87
16 $\times$, $k$=128 64.10 87.79
Table 7: Cosine similarity ($\uparrow$) between steering vectors learned by different SAE capacities. Rows/columns are ordered by expansion factor and sparsity (e × expansion, $k$ active).
SAE cfg.
16×/64
16×/512
16×/128
16×/256
8×/128
8×/512
8×/256
4×/256
4×/512
4×/128
10×/128
4×/64
8×/64
16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716
16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282
16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678
16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744
8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829
8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506
8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624
4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600
4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549
4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655
10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678
4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814
8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000
Appendix G Visual Sparse Steering Pseudocode
For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.1.
Algorithm 1 SAE_Steering – SAE-based CLS token modification
1:function SAE_Steering($\mathbf{h},\mathrm{SAE},k,\gamma,\lambda_{\Delta z},\text{mode}$) $\triangleright$ $\mathbf{h}$: CLS token; SAE: sparse autoencoder; $k$: sparsity level
2:     $\mathbf{a},\mathbf{idx}\leftarrow\mathrm{SAE.select\_topk}(\mathrm{SAE.pre\_acts}(\mathbf{h}),k)$
3:     $\mathbf{z}_{\text{base}}\leftarrow\mathrm{SAE.decode}(\mathbf{a},\mathbf{idx})$
4:     $\mathbf{z}_{\text{boost}}\leftarrow\mathrm{SAE.decode}(\gamma\cdot\mathbf{a},\mathbf{idx})$
5:     if $\text{mode}=\texttt{"reconstruction"}$" then
6:         return $\mathbf{z}_{\text{base}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
7:     else if $\text{mode}=\texttt{"amplified"}$" then
8:         return $\mathbf{z}_{\text{boost}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
9:     else if $\text{mode}=\texttt{"steering"}$" then
10:         return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$triangleright VS2
11:     end if
12:end function
Appendix H Comparison with other Baseline Methods
We compare VS2 against SpLiCE [16], using the official implementation provided by the authors. For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration. Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
Table 8: Zero-shot top-1 accuracy (%) on CIFAR-100.
Method ViT-B/32 ViT-B/16 $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) SpLiCE [16] 55.57 ( –5.50 ) 58.29 ( –5.67 ) VS2 (ours) 64.52 ( +3.45 ) 68.08* ( +4.12 )
Appendix I Effect of Top-$N$ Retrieved Neighbors
To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation. We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet. As shown in Figure 5, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$. In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples. These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
Figure 5: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors $N$, using ViT-B/16 (Patch-16) across datasets. Larger $N$ improves general classification but may degrade performance in fine-grained settings.
Appendix J Per-class Performance Analysis on CIFAR-100
In Table 9, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++. We highlight both positive and negative shifts in accuracy to identify the most impacted categories. Unlike the average performance reported in the main experiments, this analysis shows that the top-10 class-level gains are generally larger in magnitude than the corresponding losses. Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible. As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
Table 9: Top-10 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15) woman $0.57$ 0.71 $\uparrow$(+0.14) trout $0.18$ 0.32 $\uparrow$(+0.14) cockroach $0.49$ 0.62 $\uparrow$(+0.13) bridge $0.72$ 0.83 $\uparrow$(+0.11) rose $0.58$ 0.68 $\uparrow$(+0.10)
(a) VS2 – Top-10 Gains
Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07) flatfish $0.55$ 0.49 (-0.06) willow_tree $0.46$ 0.41 (-0.05) lamp $0.75$ 0.70 (-0.05) lawn_mower $1.00$ 0.96 (-0.04) kangaroo $0.72$ 0.68 (-0.04)
(b) VS2 – Top-10 Losses
Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39) raccoon $0.44$ 0.83 $\uparrow$(+0.39) lizard $0.36$ 0.74 $\uparrow$(+0.38) wolf $0.54$ 0.90 $\uparrow$(+0.36) shrew $0.44$ 0.79 $\uparrow$(+0.35) bear $0.51$ 0.85 $\uparrow$(+0.34)
(c) VS2++ – Top-10 Gains
Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03) lawn_mower $1$ 0.98 (-0.02)
(d) VS2++ – Top-10 Losses

### v2

@@
-Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors
-Abstract
-Steering vision foundation models at inference time without retraining or access to large labeled datasets is a desirable yet challenging objective, particularly in dynamic or resource-constrained settings.
-In this paper, we introduce Visual Sparse Steering (VS2), a lightweight, test-time method that guides vision models using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders without requiring contrastive data.
-Specifically, VS2 surpasses zero-shot CLIP by 4.12% on CIFAR-100, 1.08% on CUB-200, and 1.84% on Tiny-ImageNet.
-We further propose VS2++, a retrieval-augmented variant that selectively amplifies relevant sparse features using pseudo-labeled neighbors at inference time.
-With oracle positive/negative sets, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet.
-Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost.
-Finally, to better align the sparse features learned through the SAE reconstruction task with those relevant for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS).
-By incorporating a prototype-alignment loss during SAE training, using labels only during training while remaining fully test-time unsupervised, PASS consistently, though modestly, outperforms VS2, achieving a 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
+Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors
+Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors
+Gerasimos Chatzoudis    Zhuowei Li    Gemma E. Moran    Hao Wang    Dimitris N. Metaxas
+Machine Learning, ICML
+Steering vision foundation models at test time, without updating foundation-model weights or using labeled target data, is a desirable yet challenging goal.
+We present Visual Sparse Steering (VS2), a lightweight, label-free adaptation method that constructs a steering vector from sparse features extracted by a Sparse Autoencoder (SAE) trained on unlabeled in-domain training-split activations of the vision encoder.
+VS2 offers three key advantages over existing test-time adaptation methods: (1) a feature-level intervention space in sparse SAE representations; (2) efficiency, requiring only a forward pass with no test-time optimization or backpropagation; and (3) a reliability diagnostic based on SAE reconstruction loss that can skip steering when reconstruction is poor, enabling safe fallback to the baseline, a capability not standard in conventional steering vectors and test-time adaptation methods.
+Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves zero-shot top-1 accuracy by 3.45–4.12%, 0.93–1.08%, and 1.50–1.84%, respectively, while remaining forward-only and adding minimal compute overhead.
+A retrieval-based upper-bound analysis suggests substantial headroom if task-relevant sparse features can be selected reliably, motivating future work on selective feature amplification for interpretable, efficient test-time steering.
 1 Introduction
-Foundation Models (FMs) have exhibited strong generalization capabilities across a wide range of domains, from classification tasks to open-ended generation  clip ; sam ; llama2 ; llava ; gpt4 .
-Despite their impressive performance, these models largely function as black boxes, restricting their controllability, and reliability in practice.
-The standard approach to edit a model’s behavior involves first collecting a labeled dataset resembling the target objective, followed by a supervised fine-tuning (SFT) procedure upon the holistic zhang2023instruction ; dong2023abilities or partial (e.g., PEFT lester2021power ; lora ; ia3 ) parameter space.
-Nevertheless, this procedure is cumbersome and can become prohibitively expensive given the model scale.
-Consequently, Steering Vector (SV) methods turner2023activation ; zou2023representation ; liu2024incontext , as an alternative, have gained popularity in the Large Language Model (LLM) domain as they can effectively modulate the model’s generation at inference-time in a training-free manner.
-While SVs have shown promise in LLMs, their functionality in the vision domain remains unclear.
-In this work, we investigate the functionality of steering vectors in vision foundation models (VFMs).
-In particular, we use zero-shot classification tasks as a testbed to explore how SVs can be leveraged to upweight relevant visual features.
-Conventional SVs operate by first constructing a directional vector within the latent space to indicate a desired change of the functionality.
-It then linearly combines the directional vector with the original latent embeddings to craft output embeddings.
-In LLMs, SVs rely on two symbolic anchors that are inherently interpretable to determine the functional direction of the SV.
-For example, a happiness SV can be extracted using the latent embedding arithmetic corresponding to Happy - Sad.
-The formulation of SVs in LLM poses two critical challenges in vision domain.
-Given an image at inference-time: (1) Unlike symbolic language tokens, there exists tremendous redundancy in visual representations; consequently, it is unclear what features should be steered; (2) Without inherently interpretable anchors (e.g., happy, sad), towards which direction to steer is not obvious.
+Adapting vision foundation models at test time, without labeled data or weight updates, is desirable but challenging.
+Many test-time adaptation (TTA) methods rely on backpropagation through large encoders, incurring substantial compute overhead  (Shu et al., 2022; Feng et al., 2023; Yoon et al., 2024).
+Steering vectors offer a lightweight alternative but typically require carefully curated contrastive examples, limiting practicality in vision  (Turner et al., 2023; Zou et al., 2023; Liu et al., 2024).
+This motivates two challenges: (i) Can we construct effective steering directions without contrastive supervision or test-time optimization?
+and (ii) Can we find representations that mitigate the redundancy and entanglement inherent in visual features?
+Drawing inspiration from mechanistic interpretability work on sparse autoencoders (SAEs) (Elhage et al., 2022; Bricken et al., 2023), we hypothesize that sparse SAE latents capture salient aspects of the input and can provide a less redundant basis for intervention in visual representations.
+SAEs have been used primarily as an interpretability tool to extract human-interpretable features in trained foundation models.
+However, their potential as a control mechanism remains underexplored: it is unclear whether sparse features can be manipulated in a principled, label-free manner to improve downstream performance.
+This gap motivates our approach: rather than using SAE features for post-hoc analysis, we leverage them as an actionable intervention space for efficient, label-free test-time steering.
+To this end, we propose Visual Sparse Steering (VS2): we train a top-$k$ SAE on unlabeled in-domain activations and, at test time, construct a steering vector by amplifying the image’s active sparse features and projecting the induced change back to embedding space.
+VS2 is forward-only, requiring a single SAE forward pass with no test-time optimization, backpropagation, or contrastive anchor examples.
+Across CIFAR-100, CUB-200, and Tiny-ImageNet with two CLIP backbones (ViT-B/32, ViT-B/16)  (Radford et al., 2021b), VS2 improves zero-shot top-1 accuracy by 3.45%-4.12%, 0.93%-1.08%, and 1.50%-1.84%, respectively, while adding $0.1\%$ compute overhead.
+A practical challenge is distribution shift: if a test input lies outside the SAE’s training distribution, reconstruction quality can degrade, and the resulting sparse features may be unreliable, causing steering to underperform.
+To mitigate this, VS2 includes a built-in reliability diagnostic based on SAE reconstruction error.
+When reconstruction error is high, we skip steering and fall back to the original zero-shot CLIP prediction, avoiding harmful adaptation from misaligned sparse features.
+This provides a simple, model-internal mechanism for deciding when to apply steering, which is not typically available in standard steering-vector or test-time adaptation methods.
+A natural question is whether all active sparse features should be amplified uniformly, or whether selectively amplifying task-relevant features could yield larger gains.
+To quantify this headroom, we introduce VS2++ as a retrieval-based upper-bound analysis: given an external unlabeled corpus, it retrieves nearest neighbors, forms positive and negative groups using CLIP pseudo-labels, and constructs a contrastive steering direction in SAE feature space.
+With oracle positives/negatives, VS2++ yields gains of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate and motivating future work on robust selection under noisy pseudo-labels.
+In summary, our contributions are:
+  • Unsupervised sparse steering for vision.
+We introduce VS2, a lightweight, label-free method that trains a top-$k$ SAE on unlabeled in-domain activations and performs forward-only test-time steering without updating CLIP weights or using contrastive anchor examples.
+  • Consistent gains with minimal overhead.
+Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 improves zero-shot top-1 accuracy while adding $0.1\%$s compute overhead and requiring no test-time optimization.
+  • Reliability-aware fallback.
+We propose a reconstruction-error-based diagnostic that can conservatively disable steering and fall back to the baseline when the SAE representation is unreliable, reducing the risk of harmful interventions under distribution shift, at the cost of occasionally forgoing potential gains under a strict reconstruction-error threshold.
+  • Headroom via selective amplification.
+We present VS2++ as a retrieval-based upper-bound analysis showing substantial potential gains when task-relevant sparse features can be selected accurately, motivating future work on robust feature selection under noisy pseudo-labels.
 (a) VS2: Visual Sparse Steering w/o Contrastive Data
 (b) VS2++: Selective Visual Sparse Steering
-Figure 1: Overview of VS2 and VS2++.
-At inference, VS2 finds sparse latent concepts, and steers the original embedding towards the direction of amplifying those sparse features.
-When additional caching data is available, VS2 selectively enhances certain features while suppressing others in a contrastive manner.
-To address the aforementioned challenges, we propose Visual Sparse Steering (VS2).
-Drawing inspiration from the interpretability provided by sparse auto-encoders (SAEs) hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , we hypothesize that the sparse features in a SAE’s latent space can capture key aspects of a visual embedding.
-VS2 upweights these features equally to construct a steering vector for classification.
-This equal upweighting of features is imprecise in that we expect different features to be more relevant than others for classification tasks.
-Despite this, we show that the simple equal upweighting of features still improves zero-shot classification.
-To construct a more precise steering vector, we propose VS2++, which leverages additional unlabeled data to construct steering vectors for classification.
-For a given target image, VS2++ retrieves the Top-N most similar visual embeddings, and constructs positive and negative groups according to generated pseudo labels [^1] In practice, pseudo labels can be gathered using either unsupervised clustering methods or zero-shot classification results.
-The VS2++ SV is then constructed via subtracting the average SAE features of the negative group from the positive group.
-Intuitively, VS2++ upweights the SAE features most relevant for distinguishing the positive and negative groups.
-In summary, VS2 and VS2++ aim to enhance downstream classification performance while retaining the interpretability of SAEs.
-To further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS).
-VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP  radford2021learning by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively, without any external data or supervision.
-When using external unlabeled image data, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet.
-Finally, by adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
-Our main contributions are:
-  • We investigate the functionality of steering vectors in vision domain, and propose VS2 that leverages sparse feature learning to construct steering vectors for zero-shot classification.
-  • We further propose VS2++ for scenarios where additional unlabeled data is available to selectively amplify important concepts.
-In this case, we use VS2 on pseudo-positive and -negative data to upweight relevant features for classification.
-  • We propose Prototype-Aligned Sparse Steering (PASS), a lightweight method that introduces a prototype-alignment loss during SAE training using class labels only during training while remaining fully test-time unsupervised.
-  • We empirically demonstrate that VS2, VS2++, and PASS can leverage SAEs to substantially improve zero-shot classification performance, outperforming baseline methods by a large margin under their corresponding regimes, while also retaining interpretability.
+Figure 1: Overview of VS2 and VS2++: at inference time, VS2 constructs a steering vector by equally amplifying sparse features, while VS2++ selectively enhances them when an external visual corpus is available.
 2 Related Work
 Mechanistic Interpretability and Sparse Autoencoders.
-Several traditional approaches exist for interpretability in vision models, including feature visualizationsimonyan2014visualising ; zeiler2014visualizing ; olah2017feature and network dissectionbau2017network ; oikarinen2022clip .
-Mechanistic interpretability seeks to systematically analyze and understand neural networks elhage2021mathematical ; olah2020zoom , but it faces challenges due to polysemantic neurons i.e.
-units that activate in response to multiple, seemingly unrelated inputs elhage2022toy .
-This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations elhage2022toy .
-Sparse Autoencoders (SAEs) have been shown to mitigate superposition by applying sparse dictionary learning to model internals sharkey2022taking ; bricken2023monosemanticity .
-Recent efforts have leveraged SAEs to uncover interpretable features within LLMs, revealing latent units tied to grammar rules, style patterns, and factual knowledge templeton2024scaling ; cunningham2023sparse ; gao2024scaling .
-joshi2025identifiable refines this by training autoencoders on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision.
+Several traditional approaches exist for interpretability in vision models, including feature visualization (Simonyan et al., 2014; Zeiler and Fergus, 2014; Olah et al., 2017) and network dissection (Bau et al., 2017; Oikarinen and Weng, 2022).
+Mechanistic interpretability seeks to systematically analyze and understand neural networks (Elhage et al., 2021; Olah et al., 2020), but it faces challenges due to polysemantic neurons i.e.
+units that activate in response to multiple, seemingly unrelated inputs (Elhage et al., 2022).
+This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations (Elhage et al., 2022).
+Sparse Autoencoders (SAEs) have been explored to mitigate superposition by applying sparse dictionary learning to model internals (Bricken et al., 2023).
+Recent efforts have leveraged SAEs to uncover interpretable features within LLMs  (Templeton, 2024; Cunningham et al., 2023; Gao et al., 2024a).
+Joshi et al.
+(2025) train SAEs on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision.
 While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
-While some studies investigate SAEs for exploratory feature analysis or generative modeling in the visual domain hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , they primarily address concept discovery rather than performance enhancement.
-Other attempts in the vision domain try to create new SAE architectures for concept extraction  fel2025archetypal ; thasarathan2025universal .
-In contrast, our method leverages SAEs not only to interpret but also to actively steer vision models, yielding classification improvements alongside interpretable latent representations.
-Similar to our line of work, joseph2025steering validate CLIP’s steerability and its efficacy in suppressing typographic attacks.
-Patch-SAElim2024sparse improves zero-shot classification by applying latent masking based on class-level activations.
-Specifically, it aggregates SAE activations from the training split to identify the top-$k$ most frequently activated latents per class, which are then used to mask active latents at test time.
-Our approach differs in a key aspect: we do not rely on activation aggregations from the same class in training but our method is instead fully test-time unsupervised.
-This design avoids reliance on class-conditioned statistics, making our approach more flexible and applicable to a wide range of tasks without requiring dataset-specific activation distributions.
-Specifically, our method (i) operates entirely in the vision latent space, (ii) requires no class labels, external vocabularies, or gradient updates, and (iii) improves zero-shot classification by amplifying sparse SAE features rather than replacing or suppressing activations.
+While SAEs have been explored for feature analysis, generative modeling, and concept disentanglement in the visual domain (Bhalla et al., 2024; Stevens et al., 2025; Surkov et al., 2024; Fel et al., 2025; Thasarathan et al., 2025), these works focus primarily on interpretability rather than downstream performance.
+In contrast, our method leverages SAEs to actively steer vision models in a label-free, test-time setting to improve classification.
+Related efforts like (Joseph et al., 2025) study CLIP’s steerability on typographic attacks, while Patch-SAE (Lim et al., 2024) improves classification accuracy via class-conditioned latent masking.
+Our approach, on the other hand, requires no class labels and avoids class-based activation aggregation during training, enabling broader applicability without reliance on external supervision or gradient updates.
 Steering vectors.
-Steering vector (SV) methods turner2023activation ; park2023the ; hernandez2024linearity ; NIPS2013_9aa42b31 , also known as representation engineering zou2023representation , construct a directional vector (or task vector) and apply it in the latent space to change the target model’s behavior at inference-time.
-SVs have gained popularity in LLM/MLLM domain as it can improve generation security liu2024incontext , promoting truthfulness NEURIPS2023_81b83900 , mitigating hallucination vista , as well as improving inference efficiency liu2024incontext ; i2cl .
-Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts gavrikov2025can .
-In contrast, we differentiate from this line of work by focusing on Vision Foundation Models directly steering representations in the latent space without any external language signal.
-Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions bayat2025steering ; makelov2024sparse .
-Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled concept pairs, making it more adaptable to general visual representations.
-Furthermore, its applicability to vision foundation models (VFM) remains unclear.
-The proposed visual sparse steering (VS2) method demonstrates that SV methods, combined with sparse autoencoders, can be utilized to enhance image features, improving zero-shot classification.
+Steering Vector (SV) methods (Turner et al., 2023; Park et al., 2023; Hernandez et al., 2024; Mikolov et al., 2013), also known as representation engineering (Zou et al., 2023), construct a directional task vector and apply it in the latent space to change the target model’s behavior at inference time.
+In LLMs/MLLMs, SVs are used to enhance security (Liu et al., 2024), truthfulness (Li et al., 2023), reduce hallucinations (Li et al., 2025a), and improve efficiency (Li et al., 2025b).
+Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts (Gavrikov et al., 2025).
+In contrast, we focus on steering latent representations in vision models without any language input.
+Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions (Bayat et al., 2025; Makelov, 2024).
+Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled positive/negative concept pairs, making it more adaptable to general visual representations.
+Test-Time Adaptation.
+Recent advances in improving the generalization of vision-language models have introduced a variety of prompt-tuning approaches, such as CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), and MaPLe (Khattak et al., 2023), along with adapter-based techniques like Tip-Adapter (Zhang et al., 2021) and CLIP-Adapter (Gao et al., 2024b).
+Despite their effectiveness, these methods typically assume access to a small number of labeled target-domain examples (often in a few-shot setting).
+In contrast, Test-Time Adaptation methods aim to improve robustness under distribution shift by adapting a pre-trained model using only unlabeled test samples at inference time (Liu et al., 2021; Sun et al., 2020; Wang et al., 2020; Gao et al., 2022).
+Specifically, Test-Time Prompt Tuning (TPT) (Shu et al., 2022) introduced entropy-minimizing prompt optimization over augmented views, inspiring subsequent methods such as DiffTPT (Feng et al., 2023), which leverages diffusion-based augmentations, and C-TPT (Yoon et al., 2024), which incorporates calibration-aware objectives.
+Although effective, these strategies typically require optimization loops, multiple augmentations, and backpropagation through large encoders, resulting in substantial computational and memory overhead at inference.
+In contrast, VS2 performs a single forward pass with sparse SAE-guided steering, avoiding test-time optimization while still providing measurable gains.
 3 Method
-3.1 Preliminaries
-Sparse Autoencoder (SAE).
-A SAE consists of an encoder-decoder pair that maps an input vector ${\bm{x}}$ into a latent representation ${\bm{z}}$, then reconstructs ${\bm{x}}$ from ${\bm{z}}$.
-Formally, we define the encoder as:
-${\bm{z}}=\mathrm{ReLU}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}\Bigr{)},$ (1)
-where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input.
-The decoder then reconstructs ${\bm{x}}$ as:
-$\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}},$ (2)
-where ${\bm{W}}_{\mathrm{dec}}\in{\mathbb{R}}^{d\times n}$ is the decoder weight matrix.
-To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective is:
-$\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ (3)
-In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  gao2024scaling explicitly enforces the sparsity by using a top-$k$ selection mechanism:
-${\bm{z}}=\mathrm{TopK}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})\Bigr{)},$ (4)
-where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest.
+Preliminaries.
+A Sparse Autoencoder (SAE) consists of an encoder-decoder pair that maps an input vector $\bm{x}$} into a latent representation $\bm{z}$}, then reconstructs $\bm{x}$} from $\bm{z}$}.
+Formally, we define the encoder as: ${\bm{z}}=\mathrm{ReLU}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}),$ where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input.
+The decoder then reconstructs $\bm{x}$} using a decoder weight matrix , as $\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}}$ .
+To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective becomes $\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  (Gao et al., 2024a) explicitly enforces the sparsity by using a top-$k$ selection mechanism: ${\bm{z}}=\mathrm{TopK}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})),$ where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest.
 This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
 3.1 Visual Sparse Steering
-In this section, we introduce Visual Sparse Steering (VS2), a novel framework that leverages sparse features to steer neural representations toward meaningful semantic directions.
-Our approach targets two interconnected objectives, i.e., (1) explicitly identifying semantically meaningful concepts embedded within latent representations and (2) utilizing these interpretable concepts to define steering directions that enhance downstream task performance by emphasizing relevant semantic information and reducing spurious correlations.
-Unlike standard steering methods that rely on explicitly curated positive and negative examples, our approach utilizes Sparse Autoencoders (SAEs) to automatically discover and isolate sparse concepts that assist in defining a steering direction.
-By selectively amplifying these concepts within the latent space, VS2 generates conceptually grounded steering vectors that improve model interpretability, enhance classification accuracy, and robustly mitigate reliance on superficial or irrelevant cues.
+Visual Sparse Steering (VS2) is a lightweight, label-free, test-time method that guides vision models using steering vectors derived from sparse features learned by top-k Sparse Autoencoders without requiring contrastive data.
+We hypothesize that the sparse features in an SAE’s latent space can capture salient aspects of the input, thereby reducing redundancy.
+To this end, VS2 constructs a steering vector by amplifying the SAE’s salient features, reconstructing the corresponding amplified representation, and subtracting the original reconstruction (without amplification) to minimize the inherent SAE’s reconstruction errors.
+The resulting sparse steering vector points in the direction of the salient features and is linearly combined with the original embedding to steer the model’s behavior.
+Unlike common steering vectors, VS2 does not require positive and negative anchor examples; instead, steering occurs in the direction of the most salient features identified by the SAE.
 3.1.1 VS2: Visual Sparse Steering
 Learning latent concept space.
-Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding ${\bm{x}}$ into a sparse concept activation vector:
-${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|}).$ (5)
-By imposing sparsity constraints described in Sec.
-3.1 on the concept activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful concepts.
-To map the concepts into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these concepts:
-$\tilde{{\bm{x}}}=D_{c}({\bm{c}}).$ (6)
-Since explicit supervision for concept identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful concepts.
+Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding $\bm{x}$} into a sparse feature activation vector: ${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|})$.
+By imposing sparsity constraints on the feature activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful features.
+To map the features into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these features: $\tilde{{\bm{x}}}=D_{c}({\bm{c}})$.
+Since explicit supervision for feature identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful features.
 Constructing steering vector.
-At inference stage, emerged sparse concepts are deemed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$.
-In implementation, we simply scale the original concept activations ${\bm{c}}$ by a factor of $\gamma$ to amplify their effects:
-${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}.$ (7)
-Reconstructing this modified concept vector ${\bm{c}}^{\prime}$ provides a conceptually steered embedding:
-$\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime}).$ (8)
-Since reconstruction steps inherently introduce approximation errors, we define the steering vector ${\bm{v}}$ as:
-${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}},$ (9)
-which explicitly represents the semantic shift induced by upweighting and mitigates the effects of the reconstruction errors.
-Steer the target emebedding.
-Obtaining the steering vector ${\bm{v}}$ that aims to amplify the latent concepts embedded in ${\bm{x}}$, we shift the original embedding by adding the steering vector back as:
-$\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}.$ (10)
-where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale:
-$\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ (11)
-Intuitively, this procedure precisely moves the embedding ${\bm{x}}$ along directions defined by semantically relevant, interpretable concepts, thus improving interpretability and enhancing robustness against spurious correlations.
+At inference time, emerged sparse features are assumed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$.
+In implementation, we simply scale the original feature activations $\bm{c}$} by a factor of $\gamma$ to amplify their effects: ${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}$.
+Reconstructing this modified feature vector $\bm{c}$$^{\prime}$ provides a conceptually steered embedding: $\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime})$.
+Since reconstruction steps inherently introduce approximation errors, we define the steering vector $\bm{v}$} as: ${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}}$.
+which explicitly represents the semantic shift induced by upweighting and aims to mitigate the effects of the reconstruction errors.
+Steer the target embedding.
+Obtaining the steering vector $\bm{v}$} that aims to amplify the latent features embedded in $\bm{x}$}, we shift the original embedding by adding the steering vector back as: $\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}$.
+where $\lambda$ is a hyperparameter that controls the magnitude of steering.
+To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale: $\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ Intuitively, this procedure precisely moves the embedding $\bm{x}$} along directions defined by semantically relevant, salient features, thus enhancing robustness against spurious correlations.
 We refer Fig.
-1 (upper) for a schematic overview of VS2.
-3.2.2 VS2++: Visual Sparse Steering with Cached Data
-In VS2, we leverage the learned SAEs to implicitly decide the critical concepts for a given latent embedding, and treat all identified concepts equally important.
-One natural speculation is that there are certain concepts that are unique and more important to the target underlying distribution, while others are more general even spuriously correlated.
-To this end, we propose VS2++ that allows the access of cached training embedding at test-time to enable a more nuanced control over the concept steering.
+1 (left) for a schematic overview of VS2.
+3.1.2 VS2++: Visual Sparse Steering with Unlabeled Data
+In VS2, we leverage the learned SAEs to identify the most salient features for a given latent representation, but treat all identified features equally important by equally amplifying them.
+In vision models, there are potentially many redundant features, e.g., when predicting labels, we want to ignore background features such as the sky.
+To selectively amplify only important features more robustly and non-uniformly, we relax the problem, allowing the existence of an unlabeled external vision corpus, and propose VS2++ that leverages this additional unlabeled data to construct steering vectors.
+For a given target image, VS2++ retrieves the Top-N most similar visual embeddings and constructs positive and negative groups according to generated pseudo labels.
+The VS2++ Steering Vector is then constructed via subtracting the average SAE features of the negative group from the positive group.
+Intuitively, VS2++ amplifies the SAE features most relevant for distinguishing the positive and negative groups.
 Constructing Contrastive Groups.
-Given a query embedding ${\bm{x}}_{q}$ and a cached dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
+Given a query embedding ${\bm{x}}_{q}$ and an unlabeled dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
 $\mathcal{N}_{N}({\bm{x}}_{q})\;=\;\underset{\begin{subarray}{c}\{{\bm{x}}_{i}\}\subseteq\{{\bm{x}}_{i}\}_{i=1}^{M}\\ |\mathcal{N}_{N}({\bm{x}})|=N\end{subarray}}{\mathrm{arg\,max}}\;S({\bm{x}}_{q},{\bm{x}}_{i}),$ (1)
@@
 Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$.
-We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding:
-${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ (13)
-Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors:
-${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$ (14)
+We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding: ${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors: ${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$
 Selective Steering.
-For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first applies VS2 (Eq.
-7,8,9) to acquire a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying concepts within the given embedding.
+For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first obtains the steering vectors as in VS2 acquiring a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying features within the given embedding.
 We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$.
@@
 $\bar{{\bm{v}}}^{n}\;=\;\frac{1}{\bigl|{\mathbb{S}}^{-}({\bm{x}}_{q})\bigr|}\,\sum_{{\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})}{\bm{v}}_{j}^{n}.$ (3)
-Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor:
-${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ (17)
-and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
+Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor: ${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
 $\hat{{\bm{x}}}_{q}\;=\;{\bm{x}}_{q}\;+\;\lambda\,{\bm{v}},\quad\hat{{\bm{x}}}_{q}\;=\;\frac{\hat{{\bm{x}}}_{q}\,\cdot\,\|{\bm{x}}_{q}\|_{2}}{\|\hat{{\bm{x}}}_{q}\|_{2}}.$ (4)
-This selective scheme thus anchors the steering directions in concepts consistently activated within the positive group, while suppressing undesired features highlighted by the negative group.
+This selective scheme thus anchors the steering directions in features consistently activated within the positive group, while suppressing undesired features highlighted by the negative group.
 Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases.
+We refer Fig.
+1 (right) for a schematic overview of VS2++.
 4 Experiments
 Table 1: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
-CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Zero-shot (no retrieval) $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 ) RAG-enhanced (oracle cache) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle cache) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
-Experimental Setup.
-We test VS2 and VS2++ on CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny , and CUB-200 wah2011caltech three datasets, covering standard, complex, and fine-grained classification tasks.
-These datasets vary in scale, diversity, and class distribution, allowing us to assess VS2 and VS2++ in both general and imbalanced classification scenarios.
-For vision foundation model, we use CLIP radford2021learning with ViT-B/32, and ViT-B/16 backbones.
-We train top-$k$ Sparse Autoencoders (SAEs) on the CLS token embeddings extracted from every layer of CLIP’s vision transformer, using the training data of each dataset to learn sparse latent representations.
-More details regarding the training of the top-$k$ Sparse Autoencoders can be found in the Appendix C.
-4.1 Visual Sparse Steering methods
-4.1.1 VS2: Visual Sparse Steering without Contrastive Data
+CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Label-free, unsupervised steering $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( -3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( -26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( -2.85 ) 63.42 ( -0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( -20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( -17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) RAG-enhanced (oracle unlabeled external data) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle unlabeled external data) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
+Table 2: TTA comparisons.
+Method CIFAR CUB Tiny
+Zero-shot 61.07 51.76 56.64
+Ensemble 63.66 51.54 61.39
+TPT 64.09 51.83 62.77
+C-TPT 64.86 52.54 63.20
+Diff-TPT 63.04 52.80 61.60
+VS2 64.52 52.69 58.14
+VS2 + Ens 65.48 52.47 62.79
+VS2 + Ens + Aug 65.80 51.54 62.59
+Table 3: Inference-time compute.
+Method GFLOPs GMACs Params (M)
+Plain CLIP (vision) 8.7295 4.3623 87.456
+LoRA (r=16r{=}16) 8.7885 4.3918 88.046
+VS2 (SAE steering) 8.7342 4.3647 92.178
+We evaluate VS2 on three datasets: CIFAR-100 (Krizhevsky et al., 2009), Tiny-ImageNet (Le and Yang, 2015), and CUB-200 (Wah et al., 2011), covering standard, complex, and fine-grained classification tasks.
+For the vision foundation model, we use CLIP (Radford et al., 2021b) with ViT-B/32, and ViT-B/16 backbones.
+Our goal is to investigate the effectiveness of Sparse Autoencoders (SAEs) for constructing steering vectors.
+In Sections 4.1 and 4.2, we focus on the case where the SAE can faithfully reconstruct the test inputs.
+To this end, we train top-$k$ SAEs on CLS token embeddings extracted from each layer of CLIP’s vision transformer, using the training split of each dataset to learn sparse latent representations (unsupervised in-domain data).
+Additional training details are provided in Appendix C.
+In Section 4.4, we relax this assumption by training a single, more general SAE across all train datasets, including both general and fine-grained tasks, to better approximate a real-world deployment scenario and evaluate how VS2 can avoid harmful steering when test samples cannot be faithfully reconstructed.
+This capability is novel compared to common steering vector and test-time adaptation techniques.
+(a) CIFAR-100
+(b) CUB-200
+(c) Tiny-ImageNet
+Figure 2: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$.
+All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$.
+Accuracy degrades if either parameter becomes too large.
+4.1 VS2: Visual Sparse Steering
 Baselines.
-We evaluate the effectiveness of VS2 in the absence of external contrastive data.
-As baselines, we report: (1) the zero-shot performance of the original CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{CLIP}_{\text{REC}}^{\text{F}}$, which uses the SAE-reconstructed CLS token from the final layer without any steering; and (3) $\text{CLIP}_{\text{REC}}^{\text{A}}$, which uses the reconstructed CLS tokens from all layers.
-These baselines isolate the effect of reconstruction alone, without applying any modification or steering vector.
-All top-$k$ SAEs are trained on CLS token representations across all transformer layers.
-In addition to these baselines, we report $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ that applies steering by upweighting the top-$k$ sparse features of the reconstructed CLS token of the final layer with a scaling factor of $\gamma=1.5$.
-This variant isolates the impact of steering the SAE’s latent activations by amplifying their top-$k$ before the CLS embedding is reconstructed, letting us quantify the benefit attributable to this targeted up-weighting alone.
-In Appendix G, we provide the pseudocode for main differences between the proposed baselines and VS2.
-In Appendix H, we also provide comparative results of our method compared to other baselines such as Splice (bhalla2024interpreting, ).
-Note that Splice uses an external vocabulary to construct latent features; in contrast, we use unsupervised SAEs to learn latent features for VS2 and VS2++.
+We evaluate VS2 on downstream image classification and compare against baselines that isolate reconstruction and latent amplification effects.
+We report: (1) the zero-shot CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{SAE}_{\text{REC}}^{\text{F}}$, which replaces the final-layer [CLS] token with its SAE reconstruction (no steering); and (3) $\text{SAE}_{\text{REC}}^{\text{A}}$, which reconstructs [CLS] tokens from all layers (no steering).
+These baselines measure the impact of reconstruction alone.
+We additionally include $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$, which scales the top-$k$ latent activations by a fixed factor ($\gamma=1.5$) before reconstructing the final-layer [CLS] token, serving as an ablation for latent amplification without our additive steering direction.
+Appendix G provides pseudocode for these variants.
+We also report comparisons to Splice (Bhalla et al., 2024) in Appendix H, which defines latents using an external vocabulary rather than learning an unsupervised SAE dictionary.
+Unless stated otherwise, VS2 applies steering to the final-layer [CLS] token; Appendix E analyzes steering across layers.
 Results.
-From Table 1, we demonstrate that using only SAE reconstructions generally reduces zero-shot performance in all datasets; we hypothesize this is due to reconstruction errors in SAEs (engels2025decomposing, ).
-However, when SAE features are amplified in the reconstruction, performance can improve, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ for CIFAR-100.
-Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% using the ViT-B/32 and ViT-B/16 backbones respectively.
-However, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ is worse than the CLIP baseline on CUB-200 and Tiny-ImageNet.
-In contrast, our method VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively.
-In summary, we demonstrate that our proposed steering vectors using SAEs can improve downstream classification performance on vision tasks without relying on any external data or supervision.
-In Appendix A, we also provide evidence that the sparse features are interpretable; they can capture subtle class-specific attributes and concepts such as birds with gray upperpart, white underpart, white-colored throat and black eyes.
-Finally, in Appendix B, we provide a sensitivity analysis of VS2 hyperparameters $\gamma$ and $\lambda$ showing the robustness of our proposed methods.
-4.1.2 VS2++: Visual Sparse Steering with Selectively Amplifying Features
-In VS2, we leverage learned SAEs to identify critical concepts from a latent embedding, treating all activated concepts as equally important.
-However, some concepts may be more discriminative or unique to the target class, while others may be generic or even spuriously correlated with a class label.
-To enable finer control over concept importance, we introduce VS2++, which accesses cached training data at test time to guide steering more selectively.
-As a proxy for identifying salient features, we adopt a Retrieval-Augmented Generation (RAG) approach: we retrieve the top-$50$ most similar training images to the test input using DINOv2 oquab2023dinov2 and average their sparse activations.
-To construct the positive and negative sets as shown in equations 13, 14, we assign pseudolabels using a CLIP classifier; images with the most frequent pseudolabel form the positive set, while the remaining images serve as negatives.
-This setup corresponds to the Non-oracle cache in Table 1.
-For comparison, we also report an unrealistic scenario that serves as an upper-bound i.e.
-Oracle cache setting that uses ground-truth labels to define positives and negatives.
-Baselines.
-As a baseline in the Retrieval-Augmented Generation (RAG) pipeline lewis2020retrieval , we combine the test query embedding with a weighted aggregation of retrieved document embeddings.
-Additional details on the weighting strategy are provided in Appendix D.
-We also include a non-SAE variant that constructs a steering vector by averaging the CLIP embeddings of the positive and negative sets and taking their difference, yielding a purely contrastive direction without relying on Sparse Autoencoders.
+From Table 1, we show that using only SAE reconstructions typically reduces zero-shot performance; we attribute this degradation to reconstruction error in SAEs (Engels et al., 2025).
+However, when SAE features are amplified during reconstruction, performance can improve in some cases, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ on CIFAR-100.
+Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% with ViT-B/32 and ViT-B/16, respectively, but underperforms CLIP on CUB-200 and Tiny-ImageNet.
+In contrast, VS2 consistently improves over zero-shot CLIP on all three datasets and both backbones, yielding gains of 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet for ViT-B/32 and ViT-B/16, respectively.
+These results demonstrate that sparsity-guided steering outperforms both reconstruction-only and amplification-only baselines.
+Although our focus is on how to steer ViTs using SAE features rather than interpretability, SAEs are often studied in mechanistic interpretability to identify human-interpretable features.
+Accordingly, Appendix A provides qualitative examples to assess whether the learned sparse features are consistent with subtle class-specific visual attributes (e.g., gray upperparts, white underparts, white throat, and black eyes in bird images).
+Finally, Figure 2 and Appendix B report a sensitivity analysis over the steering hyperparameters $\gamma$ and $\lambda$, showing a broad region of near-optimal performance and supporting the robustness of VS2 to reasonable hyperparameter choices.
+Comparison with Test-Time Adaptation Baselines.
+Table 2 compares VS2 to test-time prompt tuning methods TPT (Shu et al., 2022), C-TPT (Yoon et al., 2024), and Diff-TPT (Feng et al., 2023) using ViT-B/32.
+These baselines adapt by optimizing text prompts at test time, whereas VS2 performs no test-time optimization and does not update prompts or model weights.
+The Ensemble setting uses 80 fixed ImageNet-style templates following CLIP’s standard zero-shot protocol (Radford et al., 2021a).
+Aug denotes test-time image augmentations.
+The TTA baselines follow the standard 63-view protocol, while our VS2+Ens+Aug ablation uses 10 augmented views per image.
+Since the compared methods use different adaptation mechanisms and test-time compute, we treat this as an accuracy reference under common TTA evaluation protocols rather than a matched-compute comparison.
+Under this evaluation, VS2 matches or exceeds prompt-tuning baselines on CIFAR-100 and remains competitive on CUB-200; moreover, combining VS2 with prompt ensembling closes much of the gap on Tiny-ImageNet (VS2+Ens: 62.79% vs.
+TPT: 62.77%).
+Computational Overhead.
+VS2 is designed to be a lightweight alternative to parameter- or prompt-based test-time adaptation.
+Table 3 reports inference-time cost for the CLIP ViT-B/32 vision encoder under different adaptation strategies.
+Adding VS2 increases vision-encoder compute by only 0.0047 GFLOPs ($0.1\%$), since the SAE operates only on the [CLS] token at a single layer.
+In contrast, a LoRA adapter with rank 16 is applied inside every transformer block, leading to higher FLOPs despite fewer trainable parameters.
+At training time, learning the SAE is also inexpensive.
+With cached activations, SAE training requires only 0.014 GFLOPs per sample, compared to roughly 26 GFLOPs per sample for LoRA or full fine-tuning of the vision encoder; full training-time FLOPs and parameter counts are reported in Appendix M.
+Finally, prompt-optimization-based TTA can be substantially more expensive at inference because it evaluates many augmented views and performs gradient-based updates.
+A single CLIP forward pass on a $224\times 224$ image with 100 text prompts costs approximately 0.59 TFLOPs (vision+text; Appendix M).
+Under the standard TPT protocol (63 augmented views and 4 optimization steps with a backward pass through the text encoder), this increases to roughly 9.8 TFLOPs per test image, i.e., $\approx 16.6\times$ the cost of plain CLIP (Appendix M).
+In contrast, VS2 adds only one SAE forward pass and keeps the per-image vision-encoder overhead below $0.1\%$ of CLIP.
+4.2 VS2++: Upper-Bound Analysis via Selective Amplification
+In VS2, we uniformly amplify the active sparse features of the test image.
+A natural extension is selective amplification: upweight features that are predictive for the target class while suppressing nuisance factors.
+To quantify the headroom available from accurate feature selection, we introduce VS2++ as a retrieval-based upper-bound analysis that uses neighbors to induce weak supervision for selecting which sparse features to steer.
+Concretely, for each test image we retrieve the top-$50$ nearest neighbors from the dataset’s unlabeled training split using DINOv2 (Oquab et al., 2023).
+We then form pseudo-positive and pseudo-negative groups using CLIP predictions: neighbors with the most frequent pseudo-label constitute positives, and the remaining neighbors are treated as negatives.
+We construct a contrastive steering direction in SAE feature space by subtracting the average negative sparse activations from the average positive sparse activations.
+This corresponds to the non-oracle setting in Table 1.
+For comparison, we also report an oracle upper bound where positives and negatives are defined using the ground-truth labels of the retrieved neighbors.
+We emphasize that VS2++ is used primarily to study the gains possible from reliable feature selection, motivating future work on robust selection under noisy pseudo-labels.
+As baselines, we include a weighted Retrieval Augmented Generation (RAG) pipeline (Lewis et al., 2020), which combines the query embedding with a weighted average of retrieved CLIP embeddings (Appendix D), as well as a non-SAE contrastive steering baseline that computes the difference between the mean CLIP embeddings of the positive and negative groups.
 Results.
-In Table 1, we present the results for VS2 ++ when external data is available under both RAG-enhanced settings where we have oracle and non-oracle cache data.
-When we have oracle positive and negative sets, we observe that VS2++ surpasses all other methods under almost every setting.
-Weighted RAG performs worse than the steering vector approaches (both CLIP and VS2++), suggesting that directionality is more informative than proximity.
-In the more realistic scenario where we do not have oracle positive and negative sets, we observe drops in accuracies.
-We hypothesize that this degradation is due to inaccuracies in pseudolabeling, resulting in the amplification of spurious directions in the SAE latent space.
-However, across the six configurations, VS2++ outperforms the CLIP Steering Vector in four cases and falls behind only in the ViT-B/32 runs for CIFAR-100 and Tiny-ImageNet (with CIFAR-100 essentially on par).
-These results suggest that there remains significant room for improvement in actively selecting which sparse features to steer, especially under noisy or weakly supervised retrieval scenarios.
-In Appendix I, we conduct ablation studies on the top-$N$ retrieved images that indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
-4.1.3 Fine-Grained Per-Class Accuracy Analysis
-To understand where Visual Sparse Steering helps most, we compute per-class top-1 accuracy on CIFAR-100 with a ViT-B/32 backbone and report the ten largest improvements in Table 2.
-The results reveal a long-tailed pattern: VS2 lifts accuracy by up to 25 % in individual classes (e.g.
-lawn-mower $\rightarrow$ tractor, pine-tree $\rightarrow$ forest), while VS2++ reaches gains of 38 % when high-quality neighbours are available (e.g.
-bee $\rightarrow$ spider, lion $\rightarrow$ tiger, whale $\rightarrow$ flatfish).
-In all cases, the corrected predictions involve visually or taxonomically proximate categories, indicating that sparse steering primarily disambiguates fine-grained confusions rather than providing a uniform boost.
-This behaviour aligns with our qualitative findings in the Appendix A.2, where the top-$k$ SAE features capture subtle class-specific attributes (birds with gray upperpart, white underpart, white-colored throat and black eyes etc.) that are otherwise entangled in the original CLIP embedding.
-A more detailed list of per-class performance accuracies of VS2 and VS2++ is provided in Table 9 in the Appendix.
-Table 2: Top-10 class gains on CIFAR-100. Numbers are top-1 accuracies; green = absolute gain over zero-shot (ZS).
-Class ZS VS2 (+$\Delta$) Mislabel Why visually confusing tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower Both are wheeled agricultural machines; similar viewed side-on.
-forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree Dense tree canopies often dominated by tall conifers.
-man $53.0$ 74.0 $\uparrow$(+21.0) boy Male human silhouette; age difference subtle at CIFAR resolution.
-bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck Rectangular vehicle profile with windows and wheels.
-snake $61.0$ 76.0 $\uparrow$(+15.0) worm Long, legless bodies with smooth textures.
-woman $57.0$ 71.0 $\uparrow$(+14.0) girl Female human figures; height/size cues are small.
-trout $18.0$ 32.0 $\uparrow$(+14.0) aquarium_fish Similar fish shape and reflective scales in water scenes.
-cockroach $49.0$ 62.0 $\uparrow$(+13.0) beetle Small, dark exoskeleton insects with segmented bodies.
-bridge $72.0$ 83.0 $\uparrow$(+11.0) sea Large spans photographed above wide blue water areas.
-rose $58.0$ 68.0 $\uparrow$(+10.0) poppy Red, multi-petal flowers with central dark core.
-(a) VS2
-Class ZS VS2++ (+$\Delta$) Mislabel Why visually confusing spider $53.0$ 91.0 $\uparrow$(+38.0) bee Small multi-legged insects; dark radial silhouettes.
-tiger $48.0$ 84.0 $\uparrow$(+36.0) lion Large orange/brown big cats; stripes vs.
-mane often unclear at low resolution.
-flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale Flattened marine shapes against blue water backgrounds.
-possum $30.0$ 65.0 $\uparrow$(+35.0) hamster Small brown furry mammals with rounded bodies.
-lizard $39.0$ 74.0 $\uparrow$(+35.0) snake Slender reptiles; legs on lizard hard to spot in $32\times 32$ images.
-sweet_pepper $58.0$ 92.0 $\uparrow$(+34.0) orange Round orange-coloured produce with glossy skins.
-raccoon $49.0$ 83.0 $\uparrow$(+34.0) skunk Mid-sized nocturnal mammals with dark fur and contrasting markings.
-shrew $46.0$ 79.0 $\uparrow$(+33.0) skunk Small ground mammals; dark elongated silhouettes blur distinguishing cues.
-snail $53.0$ 86.0 $\uparrow$(+33.0) mushroom Brown dome (shell vs.
-cap) resting on the forest floor.
-beetle $49.0$ 82.0 $\uparrow$(+33.0) cockroach Dark, shiny exoskeletal insects with similar body shapes.
-(b) VS2++
-4.2 Prototype-Aware Sparse Steering Vectors
+Table 1 reports VS2++ under oracle (upper bound) and non-oracle (CLIP pseudo-labeled) positive/negative sets.
+In the oracle setting, VS2++ achieves gains of up to 21.44% (CIFAR-100), 7.08% (CUB-200), and 20.47% (Tiny-ImageNet), demonstrating substantial headroom when feature selection is accurate.
+Weighted RAG is generally weaker than contrastive steering on CIFAR-100 and Tiny-ImageNet while remaining competitive on CUB-200, indicating that directional offsets are often more informative than proximity alone; VS2++ is competitive with CLIP-space steering but not uniformly best (e.g., Tiny-ImageNet with ViT-B/32).
+In the non-oracle setting, accuracy drops due to pseudo-label noise, yet VS2++ outperforms the CLIP steering vector in four of six configurations and does not exhibit negative transfer on CUB-200.
+Overall, these results motivate robust feature selection under weak supervision; Appendix I ablates the number of retrieved neighbors.
+Table 4: Top-5 class gains on CIFAR-100. Top-1 accuracy; green = gain over CLIP zero-shot.
+Class ZS VS2 / VS2++ Mislabel Why visually confusing VS2 tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower wheeled farm machines; side view forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree dense conifer canopy man $53.0$ 74.0 $\uparrow$(+21.0) boy age cue weak at $32{\times}32$ bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck boxy vehicle silhouette snake $61.0$ 76.0 $\uparrow$(+15.0) worm elongated legless body VS2++ spider $53.0$ 91.0 $\uparrow$(+38.0) bee small dark radial silhouette tiger $48.0$ 84.0 $\uparrow$(+36.0) lion stripes vs.
+mane unclear flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale flat marine shape on blue possum $30.0$ 65.0 $\uparrow$(+35.0) hamster small brown furry mammal lizard $39.0$ 74.0 $\uparrow$(+35.0) snake legs often invisible
+4.3 Fine-Grained Per-Class Accuracy Analysis
+[…]
+4.4 SAEs as a reliability diagnostic for safe visual sparse steering
+Our main experiments assume access to a sparse autoencoder whose reconstructions of the vision encoder activations are sufficiently faithful on the target distribution.
+We approximate this by training top-$k$ SAEs on unlabeled in-domain training-split activations for each dataset, which allows us to isolate and investigate the effectiveness of sparsity-guided steering vectors in modulating the classification behavior of CLIP’s vision transformer.
+Relaxation: a generalized SAE.
+In practice, a single high-fidelity SAE that remains faithful across diverse domains is difficult to obtain.
+To approximate this setting, we train a generalized SAE on the union of the unlabeled training splits of CIFAR-100, CUB-200, and Tiny-ImageNet and then apply this same SAE across datasets.
+Table 5 reports top-1 accuracy together with Fraction of Variance Unexplained (FVU) for generalized steering.
+Larger FVU values indicate poorer reconstruction, and FVU >1>1 implies worse reconstruction than a mean predictor.
+We observe that generalized steering improves over the zero-shot baseline on CIFAR-100 and Tiny-ImageNet, where reconstruction error is low (FVU $\approx 0.21$ on CIFAR-100 and $\approx 0.44$ on Tiny-ImageNet), but degrades on CUB-200, where reconstruction error is substantially higher (FVU $\approx 1.93$).
+This pattern suggests that steering effectiveness is sensitive to SAE fidelity, motivating the use of reconstruction error as a reliability signal.
+Table 5: Reconstruction and classification with a generalized SAE. We report top-1 accuracy (%) with corresponding reconstruction error FVU in parentheses.
+CIFAR-100 CUB-200 Tiny-IN
+Method B/32 B/16 B/32 B/16 B/32 B/16
+Baseline 61.07 (–) 63.96 (–) 51.76 (–) 55.06 (–) 56.64 (–) 61.08 (–)
+VS2 64.63 (0.216) 68.22 (0.208) 48.81 (1.931) 49.22 (1.930) 59.73 (0.448) 63.83 (0.437)
+Table 6: FVU-calibrated fallback for VS2 on in-distribution and out-of-distribution (OOD) datasets using a generalized SAE.
+Each cell reports top-1 accuracy with coverage (fraction steered) in parentheses.
+In Distribution Out of Distribution
+Method CIFAR-100 CUB-200 Tiny-IN Aircraft Food101 Flower102 Caltech101 SUN EuroSAT
+[…]
+VS2 (In Domain SAE, ungated) 64.52 52.69 58.14 27.69 81.20 64.81 80.95 54.80 34.65
+VS2 (Generalized SAE, ungated) 64.63 48.81 59.73 18.57 81.06 64.11 79.72 45.78 30.76
+VS2 + FVU gate (q=0.90q{=}0.90) 64.27 (99.93%) 51.76 (0.0%) 59.29 (85.0%) 24.87 (0.00%) 80.72 (0.00%) 63.70 (0.11%) 78.44 (2.30%) 51.80 (0.11%) 30.39 (54.20%)
+VS2 + FVU gate (q=0.95q{=}0.95) 64.62 (99.97%) 51.76 (0.04%) 58.44 (92.16%) 24.87 (0.00%) 80.71 (0.02%) 63.72 (1.32%) 78.52 (5.51%) 51.80 (0.49%) 31.39 (63.15%)
+VS2 + FVU gate (q=0.99q{=}0.99) 64.63 (100%) 51.79 (0.62%) 58.65 (98.63%) 24.87 (0.09%) 80.70 (0.89%) 64.34 (22.72%) 78.76 (23.82%) 51.92 (6.97%) 31.57 (85.52%)
+VS2 + FVU gate (q=0.995q{=}0.995) 64.63 (100%) 51.83 (1.54%) 58.64 (99.34%) 24.87 (0.21%) 80.69 (3.10%) 64.21 (40.62%) 78.78 (35.26%) 52.11 (15.28%) 31.65 (92.09%)
+Out-of-distribution steering with FVU-gated fallback.
+We next evaluate robustness under distribution shift by applying the generalized SAE to six unseen datasets: Aircraft (Maji et al., 2013), Food101 (Bossard et al., 2014), Flower102 (Nilsback and Zisserman, 2008), Caltech101 (Fei-Fei, 2004), SUN (Xiao et al., 2010), and EuroSAT (Helber et al., 2019).
+As a reference point, Table 6 shows that when an in-domain SAE is used, VS2 consistently improves over zero-shot CLIP across all six OOD datasets.
+However, ungated generalized steering can substantially degrade performance on some OOD datasets (e.g., Aircraft and SUN), consistent with unreliable sparse features when reconstruction quality is poor.
+We therefore use per-sample FVU as a reliability diagnostic: if a test sample’s FVU exceeds a threshold $\tau$, we skip steering and fall back to the baseline CLIP embedding; otherwise, we apply VS2.
+To set $\tau$ without labels, we calibrate a single threshold on the unlabeled CIFAR-100+Tiny-ImageNet training union by taking the $q$-quantile of per-sample FVU scores, and then apply the same $\tau$ across all datasets.
+For FVU-gated rows in Table 6, we report top-1 accuracy together with coverage, the fraction of samples for which steering is applied.
+Overall, FVU gating acts as a conservative safety mechanism: on datasets where generalized steering is harmful, coverage drops and performance moves toward the baseline, while on in-distribution datasets with reliable reconstruction, coverage remains high and the gains of generalized steering are largely retained.
+5 Discussion and Future Work
+We introduced VS2, a lightweight and fully label-free adaptation method that steers a frozen vision encoder using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders (SAEs).
+Unlike conventional steering vectors that rely on curated contrastive anchors, VS2 constructs a per-instance direction directly from the SAE’s sparse support: it amplifies the active latents, projects the induced change back to embedding space, and applies the resulting offset at inference time.
+Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 consistently improves zero-shot accuracy while remaining forward-only and adding negligible compute overhead.
+More broadly, our results support the view that SAE latents can serve as an actionable intervention space for improving downstream behavior, beyond post-hoc interpretability.
+A key practical constraint is SAE fidelity.
+When reconstruction is poor, the resulting sparse features can be unreliable and steering may become harmful.
+A central advantage of VS2 is that reconstruction error provides a model-internal reliability signal that is typically absent from standard steering-vector pipelines and many test-time adaptation methods.
+This enables a conservative, label-free fallback: when reconstruction error is high, we revert to the zero-shot prediction, preventing negative steering under distribution shift.
+An important future direction is to adapt the SAE per instance at test time via lightweight self-reconstruction to reduce error, improve the salient features, and therefore steering effectiveness.
+Finally, our upper-bound analyses indicate substantial headroom when task-relevant sparse features can be selected reliably.
+This highlights a central gap: the SAE objective surfaces features that are salient for reconstruction, but the features most useful for downstream classification can differ and are task-dependent.
+In our current setup, the same SAE features are used independent of the task, and we uniformly amplify the active features.
+Bridging reconstruction saliency and task saliency motivates future work on task-aligned sparse steering, i.e., selecting and amplifying the subset of SAE features that most directly support the target decision.
+Appendix K takes an initial step in this direction with Prototype-Alignment Sparse Steering, inspired by prototype theory (Rosch, 1973), suggesting that incorporating weak or indirect task structure during representation learning can better align sparse features with downstream needs.
+6 Conclusion
+We introduced Visual Sparse Steering (VS2), a lightweight, label-free domain adaptation method that steers vision foundation models using sparse features learned by top-$k$ Sparse Autoencoders, without updating model weights or requiring contrastive anchor examples.
+Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves over zero-shot CLIP by 3.45%–4.12%, 0.93%–1.08%, and 1.50%–1.84%, respectively, while remaining forward-only at inference.
+To study the potential benefits of selective feature amplification, we presented VS2++ as a retrieval-based upper-bound analysis that uses retrieved neighbors to form positive/negative groups and construct a contrastive steering direction in SAE feature space.
+With oracle positives/negatives, VS2++ yields absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate.
+When the retrieved sets are noisy, performance can degrade, highlighting the need for more robust feature selection under weak supervision.
+We additionally observe that VS2 and VS2++ can disproportionately benefit visually or taxonomically proximate classes, improving per-class accuracy by up to 25% and 38%, respectively.
+Finally, VS2 provides a reconstruction-error-based reliability diagnostic that can disable steering and fall back to the baseline under a strict threshold, reducing the risk of harmful interventions at the cost of occasionally foregoing gains.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix
+Appendix A Decoding the Sparse Latent Space: Insights from SAEs
+Our proposed methods achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs).
+We hypothesize that SAEs identify meaningful sparse features, which in turn guide the steering mechanisms.
+To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned features.
+A.1 Quantitative Evaluation of Feature Significance
+To evaluate the role of sparse features, we manipulate the top-$k$ most active latent features extracted via Sparse Autoencoders.
+This experiment examines whether these features are critical for classification and how model predictions change under different modifications.
+We explore the following manipulation settings:
+  1. 1.
+    Zeroing-Out ($\gamma=0$): We set the top-$k$ most active features to zero before applying the steering vector. This removes their influence while preserving the remaining latent structure.
+  2. 2.
+    Negation ($\gamma=-1$): We invert the sign of the top-$k$ features before applying the steering vector, effectively pushing the representation in the opposite direction.
+This tests whether these dimensions encode class-discriminative information.
+Table 7 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention.
+We observe that negating or zeroing the top-$k$ most important features consistently degrades performance across all datasets.
+This result confirms that the features learned by the Sparse Autoencoders are essential for classification.
+Table 7: Effect of manipulating top-$k$ sparse codes. Zero-shot accuracy (%) drops sharply when dominant sparse features are zeroed ($\gamma=0$) or negated ($\gamma=-1$), confirming their importance.
+Modification CIFAR-100 Tiny-IN
+$\text{CLIP}_{\text{ZS}}$ $61.07$ $56.64$
+$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $39.49$
++ Zero-out ($\gamma=0$) $1.71$ $16.11$
++ Negate ($\gamma=-1$) $0.06$ $0.82$
+A.2 Qualitative Evaluation of Features Significance
+We qualitatively investigate the features learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE).
+Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature.
+Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
+To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes.
+For the qualitative evaluation of feature significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations.
+Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images.
+Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension.
+The intuition is that if a specific concept frequently emerges among the top-activating images for a particular feature, that feature is strongly associated with that concept.
+For feature 511 as shown in Figure 3 (left), the activated images exhibit consistent semantic characteristics, including a gray upper part and a white underpart.
+Notably, this feature predominantly activates for images from similar but different classes, specifically different types of Gulls, such as Western, California, Herring, and Slaty-Backed Gulls.
+We observe that the top-activating images for these classes share all concepts in common which is not the case with dimension 3067.
+For feature 3067, as shown in Figure 3 (right) we observe that the top-activating images share common visual attributes, such as a white-colored throat and black eyes.
+Through human qualitative evaluation, we find that features in the sparse latent space capture meaningful visual concepts, grouping semantically similar images together, either from the same class or across different classes, as long as they share underlying conceptual similarities.
+Figure 3: Concept coverage analysis of learned sparse latent features in the Sparse Autoencoder (SAE).
+Each subfigure illustrates the top-activating images for two different SAE dimensions, highlighting the consistency of shared visual concepts among the highest-activated examples.
+The analysis demonstrates how sparse features capture meaningful semantic attributes, grouping semantically similar images either within or across classes.
+Left: Top-activating images for feature 511.
+The images predominantly belong to different classes of Gulls (e.g., Western, California, Herring, and Slaty-Backed Gulls), yet they share consistent visual characteristics such as a gray upper part and a white underpart.
+This suggests that feature 511 captures a semantically meaningful concept spanning multiple related categories.
+Right: Top-activating images for feature 3067.
+The images share distinct visual attributes, including a white-colored throat and black eyes.
+However, unlike feature 511, these images belong to more diverse categories, indicating that this latent dimension captures a broader concept that generalizes across different classes.
+Table 8: Top-10 most activated classes for SAE Feature 511 (left) and Feature 3067 (right).
+Feature 511 strongly aligns with gull-like classes, while Feature 3067 captures a cross-class head/throat attribute.
+Class (511) Activation Class (3067) Activation 066.Western_Gull 96.7% (29/30) 082.Ringed_Kingfisher 60.0% (18/30) 060.Glaucous_winged_Gull 96.6% (28/29) 083.White_breasted_Kingfisher 40.0% (12/30) 061.Heermann_Gull 93.3% (28/30) 079.Belted_Kingfisher 30.0% (9/30) 059.California_Gull 90.0% (27/30) 041.Scissor_tailed_Flycatcher 23.3% (7/30) 062.Herring_Gull 83.3% (25/30) 080.Green_Kingfisher 16.7% (5/30) 063.Ivory_Gull 76.7% (23/30) 002.Laysan_Albatross 13.3% (4/30) 087.Mallard 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 147.Least_Tern 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 064.Ring_billed_Gull 70.0% (21/30) 044.Frigatebird 10.0% (3/30) 084.Red_legged_Kittiwake 73.9% (17/23) 049.Boat_tailed_Grackle 10.0% (3/30)
+Class-Conditional Feature Activation.
+To provide additional evidence that SAE dimensions capture discriminative visual concepts, we report the class-conditional activation frequencies of two example features: 511 and 3067.
+Specifically, Table 8 shows the percentage of samples in the top-10 activating classes where each feature is present.
+What we observe is that feature 511 fires very often in gull-like classes.
+It also fires Larus-type seabirds, i.e., terns, kittiwakes.
+We observe that it also fires moderately on similar seabirds (Horned Puffin: 53.3%, White Pelican: 40%, Brown Pelican: 16.7%, Red-breasted Merganser: 10%).
+Additionally, we observe that Feature 511 activates near 0% on unrelated classes such as buntings, warblers, and hummingbirds.
+In contrast, Feature 3067 captures a cross-category visual attribute rather than a taxonomic grouping.
+Specifically, it consistently activates on species that share a distinctive head-and-throat pattern (e.g., white throat with darker eye region), even when those species are taxonomically unrelated.
+This indicates that the SAE possibly learns attribute-level discriminative concepts, complementing the class-specific features exemplified by Feature 511.
+Discussion.
+We note that our goal is not to claim human interpretability of SAE features.
+As shown in recent mechanistic interpretability work (Chanin et al., ), human-aligned features are often split across multiple SAE components (feature absorption), a known behavior in sparse dictionary learning.
+Nonetheless, our contribution lies in demonstrating that SAEs can go beyond interpretability, serving as effective mechanisms for performance improvement via sparse steering.
+While some features are interpretable, our main result is that sparse features enable controllable improvements in downstream classification tasks.
+Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
+We analyze how the two key hyperparameters in VS2 (i) the sparse-feature amplification $\gamma$, and (ii) the steering vector scale $\lambda$ affect downstream accuracy.
+In the absence of contrastive supervision, these parameters govern how strongly we amplify sparse activations and how far the embedding is shifted in feature space.
+We sweep both values across a grid on three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet using ViT-B/32 backbone.
+Figure 4 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees.
+Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy.
+For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$.
+Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse features and/or embedding distortion.
+The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
+(a) CIFAR-100
+(b) CUB-200
+(c) Tiny-ImageNet
+Figure 4: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$.
+All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$.
+Accuracy degrades if either parameter becomes too large.
+Appendix C Training top-$k$ Sparse Autoencoders
+We follow CLIP (Radford et al., 2021b) with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token.
+Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer.
+We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned.
+We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units.
+Training largely follows Gao et al.
+(2024a) and uses  EleutherAI (2024), with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
+We monitor reconstruction quality with the fraction of variance unexplained (FVU; lower is better), defined as
+$\text{FVU}\;=\;\frac{\lVert\mathbf{X}-\hat{\mathbf{X}}\rVert_{F}^{2}}{\lVert\mathbf{X}-\bar{\mathbf{X}}\rVert_{F}^{2}},$ (5)
+where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance).
+FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 0 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean.
+In Table 9, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
+Table 9: Fraction of variance unexplained (FVU; lower is better). Each SAE has 4× expansion; sparsity $k=64$ for ViT-B/32 and $k=256$ for ViT-B/16.
+Dataset ViT-B/32 (k=64) ViT-B/16 (k=256)
+CIFAR-100 $0.2812$ $0.1166$
+CUB-200 $0.2487$ $0.1653$
+Tiny-ImageNet $0.5060$ $0.3018$
+Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
+When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline.
+We use DINOv2 (Oquab et al., 2023) to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
+$\mathbf{E}=\alpha\mathbf{q}+(1-\alpha)\sum_{j=1}^{k}w_{j}\mathbf{r}_{j},$
+where $\mathbf{q}$ is the query embedding, $\mathbf{r}_{j}$ the $j$-th retrieved embedding, and $w_{j}$ the normalized similarity weight:
+$w_{j}=\frac{s_{j}}{\sum_{i=1}^{k}s_{i}},\quad\text{where }s_{j}=\text{sim}(\mathbf{q},\mathbf{r}_{j}).$
+The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set.
+We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance.
+Figure 5 shows results for CIFAR-100, and Tiny-ImageNet.
+CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance.
+On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result.
+Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context.
+The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines.
+For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ might work best; for noisy domains, more reliance on the retrieval images might be more beneficial.
+(a) CIFAR-100
+(b) Tiny-ImageNet
+Figure 5: RAG sensitivity to $\alpha$ and top-$k$.
+Accuracy varies with the weight $\alpha$ on the original query and the number $k$ of retrieved images.
+Larger $k$ often introduces noise; smaller $\alpha$ performs better on cluttered datasets.
+Appendix E Steering Across Layers of the CLS Token
+Our default method reconstructs and applies steering to the CLS token embedding at the final layer of the vision encoder.
+To study depth effects, we evaluate steering the CLS token across multiple final layers of the transformer.
+We use ViT-B/16 on CIFAR-100 with expansion factor 4 and 128 sparse activations.
+Table 10 reports top-1 accuracy when steering the last 1, 2, or 3 layers.
+Steering only the final layer yields the best performance.
+Adding earlier layers causes a progressive drop in accuracy, falling below the zero-shot baseline when steering the last three layers.
+Table 10: Effect of steering depth on CLS. Top-1 accuracy (%) on CIFAR-100 (ViT-B/16) as the number of steered final layers increases.
+Steered Layers Accuracy (%)
+12 68.08
+11 + 12 65.72
+10 + 11 + 12 59.36
+Applying steering at multiple layers likely introduces compounding perturbations that propagate forward, making later representations harder to align with the classifier.
+CLS steering is most effective at the final layer.
+Future work could explore coordinated multi-layer steering to avoid error accumulation.
+Appendix F Ablation Study of Expansion factor and top-$k$
+In Table 11, we present the downstream task accuracy of CLIP ViT-B/32 using various values of expansion factor and $k$.
+Across a 4 $\times$ range in width and an 8 $\times$ range in sparsity, top-1 accuracy fluctuates by less than one percentage point evidence that VS2’s performance is largely insensitive to the precise SAE capacity–sparsity trade-off.
+Additionally, in Table 18, we present the average cosine similarities of different steering vectors coming from various configurations of SAEs in terms of expansion factor and top-$k$.
+Table 11: VS2 accuracy as a function of SAE width (expansion factor) and sparsity (top-$k$).
+Numbers are top-1 / top-5 (%).
+The best result is boldfaced; every other configuration is within of the optimum, highlighting the method’s robustness to architectural choices.
+SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
+4 $\times$, kk=128 64.61 87.95
+8 $\times$, kk=128 64.56 87.76
+4 $\times$, kk=64 64.54 87.79
+16 $\times$, kk=64 64.54 87.78
+8 $\times$, kk=64 64.52 87.96
+10 $\times$, kk=128 64.43 87.81
+16 $\times$, kk=512 64.42 87.71
+8 $\times$, kk=512 64.40 87.91
+4 $\times$, kk=256 64.34 87.62
+8 $\times$, kk=256 64.29 87.80
+16 $\times$, kk=256 64.28 87.75
+4 $\times$, kk=512 64.12 87.87
+16 $\times$, kk=128 64.10 87.79
+Appendix G Visual Sparse Steering Pseudocode
+For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.
+Algorithm 1 SAE_Steering – SAE-based CLS token modification
+1:function SAE_Steering($\mathbf{h},\mathrm{SAE},k,\gamma,\lambda_{\Delta z},\text{mode}$) $\triangleright$ $\mathbf{h}$: CLS token; SAE: sparse autoencoder; $k$: sparsity level
+2:  $\mathbf{a},\mathbf{idx}\leftarrow\mathrm{SAE.select\_topk}(\mathrm{SAE.pre\_acts}(\mathbf{h}),k)$
+3:  $\mathbf{z}_{\text{base}}\leftarrow\mathrm{SAE.decode}(\mathbf{a},\mathbf{idx})$
+4:  $\mathbf{z}_{\text{boost}}\leftarrow\mathrm{SAE.decode}(\gamma\cdot\mathbf{a},\mathbf{idx})$
+5:  if $\text{mode}=\texttt{"reconstruction"}$ then
+6:   return $\mathbf{z}_{\text{base}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
+7:  else if $\text{mode}=\texttt{"amplified"}$ then
+8:   return $\mathbf{z}_{\text{boost}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
+9:  else if $\text{mode}=\texttt{"steering"}$ then
+10:   return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$ $\triangleright$ VS2
+11:  end if
+12:end function
+Appendix H Comparison with other Baseline Methods
+We compare VS2 against SpLiCE (Bhalla et al., 2024), using the official implementation provided by the authors.
+For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration.
+Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
+Table 12: Zero-shot top-1 accuracy (%) on CIFAR-100.
+Method ViT-B/32 ViT-B/16
+$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 )
+$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 )
+$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 )
+$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 )
+SpLiCE (Bhalla et al., 2024) 55.57 ( –5.50 ) 58.29 ( –5.67 )
+VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 )
+Appendix I Effect of Top-$N$ Retrieved Neighbors
+To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation.
+We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet.
+As shown in Figure 6, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$.
+In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples.
+These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks may benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
+Figure 6: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors NN, using ViT-B/16 (Patch-16) across datasets.
+Larger NN improves general classification but may degrade performance in fine-grained settings.
+Appendix J Per-class Performance Analysis on CIFAR-100
+In Table 17, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++.
+We highlight both positive and negative shifts in accuracy to identify the most impacted categories.
+Unlike the average performance reported in the main experiments, this analysis shows that the top-5 class-level gains are generally larger in magnitude than the corresponding losses.
+Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible.
+As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
+Appendix K Prototype-Aware Sparse Steering Vectors
 Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE).
-These directions correspond to latent concepts that, ideally, align with class-discriminative features.
+These directions correspond to latent features that, ideally, align with class-discriminative concepts.
 Steering in these directions upweights what the model has learned to be important during reconstruction.
@@
 In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task.
-In this section, drawing inspiration from prototype theory rosch1973natural , we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
+In this section, drawing inspiration from prototype theory (Rosch, 1973), we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
 The limited improvements observed on fine-grained datasets like CUB-200 and Tiny-ImageNet suggest that the challenge lies not just in identifying sparse features, but in uncovering the correct ones.
 This shifts the central question from “What are the most important sparse features to select?” to a deeper inquiry: “Can the reconstruction objective alone reliably capture features that are most useful for classification and if not, how can task-relevant information be effectively incorporated?”
+Table 13: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and Without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
+CIFAR-100 CUB-200 Tiny-IN
+Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16
+Zero-shot (no retrieval)
+$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 )
+$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 )
+$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 )
+$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 )
+VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 )
+VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 )
 Oracle steering with known prototypes.
@@
 This confirms the existence of discriminative sparse directions.
-In Appendix E we further examine the discriminative ability (measured by orthogonality) of these steering vectors.
+In Appendix L we further examine the discriminative ability (measured by orthogonality) of these steering vectors.
 Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
@@
 Gains are modest on CUB-200 and Tiny-ImageNet, and we thus hypothesize that classes which share many features require richer or multi-prototype guidance which is an intriguing avenue for future work.
-5 Conclusion
-We introduce Visual Sparse Steering VS2 and its retrieval-augmented variant VS2++, two test-time methods that steer vision foundation models using sparse features obtained from top-$k$ Sparse Autoencoders without requiring contrastive data.
-Specifically, VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively.
-To control concept importance more precisely, we propose VS2++.
-With oracle positive/negative sets, VS2++ surpasses every baseline in nearly every configuration, achieving absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet.
-When those sets are noisy, its accuracy drops but still surpasses the standard Steering Vector in most cases, highlighting the need for better feature selection under weakly supervised retrieval.
-Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost.
-Finally, to further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we proposed Prototype-Aligned Sparse Steering (PASS).
-By adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
-Selecting appropriate prototypes or exploring multi-prototype guidance is an intriguing avenue for future work.
-Appendix
-Appendix A Decoding the Sparse Latent Space: Insights from SAEs
-Our proposed methods, VS2 and VS2++, achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs).
-We hypothesize that SAEs identify meaningful sparse concepts, which in turn guide the steering mechanisms.
-To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned concepts.
-A.1 Quantitative Evaluation of Concept Significance.
-To further evaluate the role of interpretable sparse features, we manipulate the top-$k$ most active latent dimensions extracted via Sparse Autoencoders.
-This experiment examines whether these features are critical for classification and how model predictions change under different modifications.
-We explore the following manipulation settings:
-  1. 1.
-    Zeroing-Out ($\gamma=0$): We set the top-$k$ most active features to zero before applying the steering vector. This removes their influence while preserving the remaining latent structure.
-  2. 2.
-    Negation ($\gamma=-1$): We invert the sign of the top-$k$ features before applying the steering vector, effectively pushing the representation in the opposite direction.
-This tests whether these dimensions encode class-discriminative information.
-Table 3 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention.
-We observe that negating or zeroing the top-$k$ most important concepts consistently degrades performance across all datasets.
-This result confirms that the concepts learned by the Sparse Autoencoders are essential for classification.
-Table 3: Effect of manipulating top-$k$ sparse codes. Zero-shot accuracy (%) drops sharply when dominant sparse features are zeroed ($\gamma=0$) or negated ($\gamma=-1$), confirming their importance.
-Modification CIFAR-100 CUB-200 Tiny-IN
-$\text{CLIP}_{\text{ZS}}$ $61.07$ $38.68$ $56.64$
-$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $37.30$ $39.49$
-+ Zero-out ($\gamma=0$) $1.71$ $1.14$ $16.11$
-+ Negate ($\gamma=-1$) $0.06$ $0.00$ $0.82$
-A.2 Qualitative Evaluation of Concept Significance.
-We qualitatively investigate the concepts learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE).
-Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature.
-Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
-To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes.
-For the qualitative evaluation of concept significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations.
-Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images.
-Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension.
-The intuition is that if a specific concept frequently emerges among the top-activating images for a particular dimension, that dimension is strongly associated with that concept.
-For feature 511 as shown in Figure 2 (left), the activated images exhibit consistent semantic characteristics, including a gray upper part and a white underpart.
-Notably, this feature predominantly activates for images from similar but different classes, specifically different types of Gulls, such as Western, California, Herring, and Slaty-Backed Gulls.
-We observe that the top-activating images for these classes share all concepts in common which is not the case with dimension 3067.
-For feature 3067, as shown in Figure 2 (right) we observe that the top-activating images share common visual attributes, such as a white-colored throat and black eyes.
-Through human qualitative evaluation, we find that features in the sparse latent space capture meaningful visual concepts, grouping semantically similar images together, either from the same class or across different classes, as long as they share underlying conceptual similarities.
-Figure 2: Concept coverage analysis of learned sparse latent features in the Sparse Autoencoder (SAE).
-Each subfigure illustrates the top-activating images for two different SAE dimensions, highlighting the consistency of shared visual concepts among the highest-activated examples.
-The analysis demonstrates how sparse features capture meaningful semantic attributes, grouping semantically similar images either within or across classes.
-Left: Top-activating images for feature 511.
-The images predominantly belong to different classes of Gulls (e.g., Western, California, Herring, and Slaty-Backed Gulls), yet they share consistent visual characteristics such as a gray upper part and a white underpart.
-This suggests that feature 511 captures a semantically meaningful concept spanning multiple related categories.
-Right: Top-activating images for feature 3067.
-The images share distinct visual attributes, including a white-colored throat and black eyes.
-However, unlike feature 511, these images belong to more diverse categories, indicating that this latent dimension captures a broader concept that generalizes across different classes.
-Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
-We analyze how the two key hyperparameters in VS2 (i) the sparse-feature amplification $\gamma$, and (ii) the steering vector scale $\lambda$ affect downstream accuracy.
-In the absence of contrastive supervision, these parameters govern how strongly we amplify sparse activations and how far the embedding is shifted in feature space.
-We sweep both values across a grid on three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet using ViT-B/32 backbone.
-Figure 3 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees.
-Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy.
-For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$.
-Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse codes and embedding distortion.
-The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
-(a) CIFAR-100
-(b) CUB-200
-(c) Tiny-ImageNet
-Figure 3: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$.
-All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$.
-Accuracy degrades if either parameter becomes too large.
-Appendix C Training top-$k$ Sparse Autoencoders
-We follow CLIP [18] with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token.
-Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer.
-We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned.
-We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units.
-Training largely follows [31] and uses  [53], with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
-We monitor reconstruction quality with the fraction of variance unexplained (FVU; lower is better), defined as
-$\text{FVU}\;=\;\frac{\lVert\mathbf{X}-\hat{\mathbf{X}}\rVert_{F}^{2}}{\lVert\mathbf{X}-\bar{\mathbf{X}}\rVert_{F}^{2}},$ (20)
-where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance).
-FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 00 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean.
-In Table 4, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
-Table 4: Fraction of variance unexplained (FVU; lower is better). Each SAE has 4× expansion; sparsity $k=64$ for ViT-B/32 and $k=256$ for ViT-B/16.
-Dataset ViT-B/32 (k=64) ViT-B/16 (k=256)
-CIFAR-100 $0.2812$ $0.1166$
-CUB-200 $0.2487$ $0.1653$
-Tiny-ImageNet $0.5060$ $0.3018$
-Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
-When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline.
-We use DINOv2 [50] to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
-$\mathbf{E}=\alpha\mathbf{q}+(1-\alpha)\sum_{j=1}^{k}w_{j}\mathbf{r}_{j},$
-where $\mathbf{q}$ is the query embedding, $\mathbf{r}_{j}$ the $j$-th retrieved embedding, and $w_{j}$ the normalized similarity weight:
-$w_{j}=\frac{s_{j}}{\sum_{i=1}^{k}s_{i}},\quad\text{where }s_{j}=\text{sim}(\mathbf{q},\mathbf{r}_{j}).$
-The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set.
-We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 4 shows results for CIFAR-100, and Tiny-ImageNet.
-Findings.
-CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance.
-On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result.
-This suggests that the retrieval set carries strong distributional signal in cluttered datasets.
-Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context.
-The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines.
-For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ work best; for noisy domains, more reliance on the retrieval pool is beneficial.
-(a) CIFAR-100
-(b) Tiny-ImageNet
-Figure 4: RAG sensitivity to $\alpha$ and top-$k$.
-Accuracy varies with the weight $\alpha$ on the original query and the number $k$ of retrieved images.
-Larger $k$ often introduces noise; smaller $\alpha$ performs better on cluttered datasets.
-Appendix E Are exemplar-derived directions really distinct?
+Table 14: Trade-off between reconstruction fidelity and prototype alignment. Increasing $w_{\text{aux}}$ improves classification accuracy but degrades FVU reconstruction.
+$w_{\text{aux}}$ FVU $\downarrow$ Accuracy (%) $\uparrow$
+0.1 0.3437 68.80
+0.5 0.4887 70.40
+1.0 0.5393 70.61
+2.0 0.5702 70.76
+Reconstruction vs. Alignment Trade-off
+Sparse Autoencoders (SAEs) trained for reconstruction can also be optimized to align their latent features with class-level prototypes.
+However, this introduces a trade-off between two competing objectives: fidelity of reconstruction and discriminative alignment.
+To investigate this trade-off, we introduce a weighting coefficient $w_{\text{aux}}$ that controls the strength of prototype alignment relative to the reconstruction loss.
+As $w_{\text{aux}}$ increases, alignment is encouraged more strongly.
+Table 14 reports the resulting changes in reconstruction loss (measured by FVU) and top-1 classification accuracy on CIFAR-100 using ViT-B/16, with 128 sparse latents and expansion factor 4.
+We observe that increasing $w_{\text{aux}}$ consistently improves classification performance, up to a point, even though it introduces more reconstruction error.
+This aligns with our hypothesis: while exact input reconstruction encourages general feature coverage, alignment with class prototypes promotes discriminative feature extraction.
+These results highlight the flexibility of SAE-based steering to balance interpretability and performance depending on downstream objectives.
+Appendix L Are exemplar-derived directions really distinct?
 A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B.
-Using the oracle prototypes as described in Section 4.2, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder.
-Most pairs have low similarity ($\!\text{mean}\!=\!0.23$), yet a non-negligible tail reveals strongly overlapping directions.
-Table 5 lists the ten highest-overlap pairs:
+Using the oracle prototypes as described in Appendix K, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder.
+Most pairs have low similarity (mean=0.23), yet a non-negligible tail reveals strongly overlapping directions.
+Table 15 lists the ten highest-overlap pairs.
 Table 15: Top-10 most overlapping prototype steering directions on CIFAR-100.
@@
 10 bowl cup $0.80$
+Table 16: Training-time compute and parameter cost per sample for various adaptation strategies on the CLIP ViT-B/32 vision encoder.
+Method Train FLOPs / sample (fwd+bwd) GFLOPs Trainable Params
+SAE (cached activations) 14,164,992 0.014 4,722,432
+SAE (non-cached; incl. extraction) 8,743,645,440 8.744 4,722,432
+LoRA (rank = 16) 26,365,388,544 26.365 589,824
+Full fine-tuning (vision encoder) 26,188,441,344 26.188 87,456,000
 These high-overlap pairs are semantically plausible confusions (e.g.
@@
 Investigating principled ways to encourage orthogonality during SAE training is a promising direction for future work.
-Appendix F Ablation Study of Expansion factor and top-$k$
-In Table 6, we present the downstream task accuracy of CLIP ViT-B/32 using various values of expansion factor and $k$.
-Across a 4 $\times$ range in width and an 8 $\times$ range in sparsity, top-1 accuracy fluctuates by less than one percentage point evidence that VS2’s performance is largely insensitive to the precise SAE capacity–sparsity trade-off.
-Additionally, in Table 7, we present the average cosine similarities of different steering vectors coming from various configurations of SAEs in terms of expansion factor and top-$k$.
-Table 6: VS2 accuracy as a function of SAE width (expansion factor) and sparsity (top-$k$).
-Numbers are top-1 / top-5 (%).
-The best result is boldfaced; every other configuration is within of the optimum, highlighting the method’s robustness to architectural choices.
-SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
-4 $\times$, $k$=128 64.61 87.95
-8 $\times$, $k$=128 64.56 87.76
-4 $\times$, $k$=64 64.54 87.79
-16 $\times$, $k$=64 64.54 87.78
-8 $\times$, $k$=64 64.52 87.96
-10 $\times$, $k$=128 64.43 87.81
-16 $\times$, $k$=512 64.42 87.71
-8 $\times$, $k$=512 64.40 87.91
-4 $\times$, $k$=256 64.34 87.62
-8 $\times$, $k$=256 64.29 87.80
-16 $\times$, $k$=256 64.28 87.75
-4 $\times$, $k$=512 64.12 87.87
-16 $\times$, $k$=128 64.10 87.79
+Table 17: Top-5 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
+Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15)
+(a) VS2 – Top-5 Gains
+Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07)
+(b) VS2 – Top-5 Losses
+Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39)
+(c) VS2++ – Top-5 Gains
+Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03)
+(d) VS2++ – Top-5 Losses
+Appendix M Computational Overhead: Additional Details
+Training-time FLOPs.
+To quantify the cost of training the SAE versus alternative adaptation methods, Table 16 reports per-sample training FLOPs (forward + backward) and the number of trainable parameters for the CLIP ViT-B/32 vision encoder.
+Even when activation extraction is included, SAE training remains more than $3\times$ cheaper than LoRA or full fine-tuning.
+With cached activations, as in our VS2 setup, the cost drops to just 0.014 GFLOPs per sample.
+TPT FLOP accounting.
+For completeness, we outline how the $16.6\times$ test-time overhead factor for TPT (Shu et al., 2022) is obtained.
+A single CLIP ViT-B/32 forward pass on a $224\times 224$ image with 100 text prompts costs $8.7$ GFLOPs for the vision encoder and $581.8$ GFLOPs for the text encoder, for a total of approximately $0.59$ TFLOPs.
+In our implementation, TPT performs $63$ augmented views per image and runs $4$ optimization steps; each step requires a full CLIP forward pass over all views and a backward pass through the text encoder.
+This yields about $9.217$ TFLOPs of adaptation compute per test image, followed by one final forward-only evaluation of $0.591$ TFLOPs, for a total of roughly $9.807$ TFLOPs.
+Dividing by the $0.59$ TFLOPs of plain CLIP gives the reported $\approx 16.6\times$ test-time compute overhead.
 Table 18: Cosine similarity ($\uparrow$) between steering vectors learned by different SAE capacities. Rows/columns are ordered by expansion factor and sparsity (e × expansion, $k$ active).
 SAE cfg.
-16×/64
-16×/512
-16×/128
-16×/256
-8×/128
-8×/512
-8×/256
-4×/256
-4×/512
-4×/128
-10×/128
-4×/64
-8×/64
-16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716
-16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282
-16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678
-16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744
-8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829
-8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506
-8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624
-4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600
-4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549
-4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655
-10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678
-4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814
-8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000
-Appendix G Visual Sparse Steering Pseudocode
-For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.1.
-Algorithm 1 SAE_Steering – SAE-based CLS token modification
-1:function SAE_Steering($\mathbf{h},\mathrm{SAE},k,\gamma,\lambda_{\Delta z},\text{mode}$) $\triangleright$ $\mathbf{h}$: CLS token; SAE: sparse autoencoder; $k$: sparsity level
-2:     $\mathbf{a},\mathbf{idx}\leftarrow\mathrm{SAE.select\_topk}(\mathrm{SAE.pre\_acts}(\mathbf{h}),k)$
-3:     $\mathbf{z}_{\text{base}}\leftarrow\mathrm{SAE.decode}(\mathbf{a},\mathbf{idx})$
-4:     $\mathbf{z}_{\text{boost}}\leftarrow\mathrm{SAE.decode}(\gamma\cdot\mathbf{a},\mathbf{idx})$
-5:     if $\text{mode}=\texttt{"reconstruction"}$" then
-6:         return $\mathbf{z}_{\text{base}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
-7:     else if $\text{mode}=\texttt{"amplified"}$" then
-8:         return $\mathbf{z}_{\text{boost}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
-9:     else if $\text{mode}=\texttt{"steering"}$" then
-10:         return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$triangleright VS2
-11:     end if
-12:end function
-Appendix H Comparison with other Baseline Methods
-We compare VS2 against SpLiCE [16], using the official implementation provided by the authors.
-For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration.
-Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
-Table 8: Zero-shot top-1 accuracy (%) on CIFAR-100.
-Method ViT-B/32 ViT-B/16 $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) SpLiCE [16] 55.57 ( –5.50 ) 58.29 ( –5.67 ) VS2 (ours) 64.52 ( +3.45 ) 68.08* ( +4.12 )
-Appendix I Effect of Top-$N$ Retrieved Neighbors
-To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation.
-We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet.
-As shown in Figure 5, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$.
-In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples.
-These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
-Figure 5: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors $N$, using ViT-B/16 (Patch-16) across datasets.
-Larger $N$ improves general classification but may degrade performance in fine-grained settings.
-Appendix J Per-class Performance Analysis on CIFAR-100
-In Table 9, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++.
-We highlight both positive and negative shifts in accuracy to identify the most impacted categories.
-Unlike the average performance reported in the main experiments, this analysis shows that the top-10 class-level gains are generally larger in magnitude than the corresponding losses.
-Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible.
-As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
-Table 9: Top-10 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
-Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15) woman $0.57$ 0.71 $\uparrow$(+0.14) trout $0.18$ 0.32 $\uparrow$(+0.14) cockroach $0.49$ 0.62 $\uparrow$(+0.13) bridge $0.72$ 0.83 $\uparrow$(+0.11) rose $0.58$ 0.68 $\uparrow$(+0.10)
-(a) VS2 – Top-10 Gains
-Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07) flatfish $0.55$ 0.49 (-0.06) willow_tree $0.46$ 0.41 (-0.05) lamp $0.75$ 0.70 (-0.05) lawn_mower $1.00$ 0.96 (-0.04) kangaroo $0.72$ 0.68 (-0.04)
-(b) VS2 – Top-10 Losses
-Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39) raccoon $0.44$ 0.83 $\uparrow$(+0.39) lizard $0.36$ 0.74 $\uparrow$(+0.38) wolf $0.54$ 0.90 $\uparrow$(+0.36) shrew $0.44$ 0.79 $\uparrow$(+0.35) bear $0.51$ 0.85 $\uparrow$(+0.34)
-(c) VS2++ – Top-10 Gains
-Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03) lawn_mower $1$ 0.98 (-0.02)
-(d) VS2++ – Top-10 Losses
+16×/64 16×/512 16×/128 16×/256 8×/128 8×/512 8×/256 4×/256 4×/512 4×/128 10×/128 4×/64 8×/64 16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716 16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282 16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678 16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744 8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829 8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506 8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624 4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600 4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549 4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655 10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678 4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814 8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000